### v1

# Price of Uncertainty for Consensus Games

Abstract
Many game-theoretic models assume that players have access to accurate information, but uncertainty in observed data is frequently present in real-world settings. In this paper, we consider a model of uncertainty where adversarial perturbations of relative magnitude $1+\varepsilon$ are introduced to players’ observed costs. The effect of uncertainty on social cost is referred to as the price of uncertainty. We prove a tight bound on the price of uncertainty for consensus games of $\Theta(\varepsilon^{2}n^{2})$ for all $\varepsilon=\Omega\mathopen{}\left(n^{-1/4}\right)$. This improves a previous lower bound of $\Omega(\varepsilon^{3}n^{2})$ as well as a previous upper bound of $O(\varepsilon n^{2})$.
1 Introduction
If you are asked whether more of your friends like vanilla ice cream or chocolate ice cream, could you truly give the right answer with confidence? Even the best polling organizations sometimes fail to predict the outcomes of two-party majority elections. The truth is that uncertainty is everywhere and can impact our abilities to make rational decisions. However, most game-theoretic models assume that players have access to accurate information about interactions as well as their own costs. In the context of strategic interactions, it is interesting to see how uncertainty can affect agents’ behavior.
Balcan et al. (2009) consider how introducing small perturbations in observed costs in well-studied potential games can impact their behavior under improved-response and best-response dynamics. They coin the term price of uncertainty, and study it in a wide variety of potential games including set-covering games (Buchbinder et al., 2008), fair cost-sharing games (Anshelevich et al., 2008), job-scheduling games (Nisan et al., 2007), and consensus games. In some games, even extremely small perturbations can cause these dynamics to snowball out of control and move to states of much higher social cost, whereas in other games these dynamics are much more stable against large degrees of perturbation.
Consensus games are a special case of party affiliation games, which are in turn a class of potential games (Bhalgat et al., 2010). In a consensus game, each of $n$ players can choose between two parties: the white group and the red group. Players accrue a constant positive cost for interacting with those that chose the other side. Players are modeled as vertices of an undirected graph that can be colored either white or red, and a player’s cost is the number of neighbors with the other color, which we call bad neighbors. We will also call edges whose endpoints have different colors bad edges.
Starting from a ground state with social cost at least 1, the price of uncertainty is the maximum multiplicative factor by which the social cost can increase via uncertain best-response dynamics. The term uncertain here refers to the fact that a player can miscount their cost, which recall is the number of bad neighbors, by a multiplicative factor of at most $1+\varepsilon$ for a small uncertainty factor $\varepsilon>0$ that can depend on $n$. Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise a player cannot miscount its number of bad neighbors and hence there is no uncertainty. Balcan et al. (2011) prove an upper bound of $O(\varepsilon n^{2})$ on the price of uncertainty for consensus games that holds for all $\varepsilon=\Omega\left(\frac{1}{n}\right)$ as well as a lower bound of $\Omega(\varepsilon^{3}n^{2})$ that holds in the domain $\varepsilon=\Omega\left(n^{-1/3}\right)$.
This paper obtains a lower bound for the price of uncertainty of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$. This is achieved using a construction that amplifies the snowball effect of uncertainty described by Balcan et al. (2011). Then, we will prove that the price of uncertainty for consensus games is also bounded by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega\left(\varepsilon^{-1/4}\right)$ through the construction and analysis of a clever monovariant. These bounds coincide, so we conclude that the price of uncertainty for consensus games is $\Theta\left(\varepsilon^{2}n^{2}\right)$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$, resolving an open problem proposed by Balcan et al. (2011).
2 Lower Bound
We first improve the lower bound for the price of uncertainty for consensus games. Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
Theorem 1.
The price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
To prove their lower bound, Balcan et al. (2009) used a carefully tuned gadget to create a snowball-like effect of uncertainty. To prove Theorem 1, we present a more complicated construction that amplifies the snowball effect even further.
2.1 Proof of Improved Lower Bound
Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty. By definition, an uncertain best response must involve a color switch. One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level. Starting from the first level, there are enough bad edges for all vertices in the first level to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$. The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
We now make the above idea rigorous. we define a boosting operation between layers of the following form. Let $\ell>1$ and suppose that level $\ell-1$ has $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m$ vertices for some $m$, level $\ell$ has $m$ vertices, and level $\ell+1$ has $\left(1+\frac{\varepsilon}{2}\right)\cdot m$ vertices. Suppose that every vertex in level $\ell$ is connected to every vertex in level $\ell-1$ and level $\ell+1$, so the degree of every vertex in level $\ell$ is $d_{\ell}=\left(\frac{1}{1+\frac{\varepsilon}{2}}+(1+\frac{\varepsilon}{2})\right)\cdot m$. If the vertices in levels $\ell$ and $\ell+1$ are all white and the vertices in level $\ell-1$ are all red, then the edges between levels $\ell-1$ and $\ell$ are all bad. Since $(1+\varepsilon)\cdot\frac{1}{1+\frac{\varepsilon}{2}}\geq\frac{d_{\ell}}{2}$ for all $\varepsilon\leq 1$ then all vertices in level $\ell$ are allowed to play uncertain best responses in sequence, turning level $\ell$ all red. Note now that the number of bad edges between levels $\ell$ and $\ell+1$ is $\left(1+\frac{\varepsilon}{2}\right)\cdot m^{2}$, which is a $1+O(\varepsilon)$ multiplicative increase from the number of original bad edges $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m^{2}$ from before this boosting operation.
We now present the construction. In order to implement the idea above, we must first find some way to create the first two layers of size $\frac{1}{\varepsilon}$ since a vertex must have degree at least $\frac{1}{\varepsilon}$ in order to play an uncertain best response. The initializer layers will have numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{1}{\varepsilon},\frac{1}{\varepsilon}$
where the odd indexed layers cycle the consecutive integers from $\frac{1}{\varepsilon}$ to $\frac{2}{\varepsilon}$ and the numbers of vertices in even indexed layers increase by a factor of 2 whenever the odd indexed layer’s cycle repeats until the number reaches $\frac{1}{\varepsilon}$. Each vertex is connected to all vertices in the layer immediately after it except for vertices in odd indexed layers with $\frac{2}{\varepsilon}$ vertices, which are connected to the next layer to form a regular bipartite graph with the next layer where the vertices in the next layer are each connected to $\frac{1}{\varepsilon}$ vertices in the current layer. Note that the total number of initializer layers is $\frac{1}{\varepsilon}\cdot\log_{2}\frac{1}{\varepsilon}$ and each layer has at most $\frac{2}{\varepsilon}$ vertices, so the total number of vertices in the initializer layers is $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$. The initializer sequence has been constructed in such a way that that starting from a state in which the first layer is red and all other layers are white, the vertices in each layer can play uncertain best responses to change their color to red in sequence so that the last two layers form a $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ where the left side is red and the right half is white. Finally, we construct $\frac{1}{\varepsilon}$ white vertices in a parallel layer to the red left side of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ and connect each of them to every vertex in the right half of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$, so that each vertex in the right half now has $\frac{1}{\varepsilon}$ red neighbors and $\frac{1}{\varepsilon}$ white neighbors on the left. Note that no new bad edges are added. We will now leave all layers before the current layer frozen, that is, no layer in a layer before the current layer will ever change color again.
After the current layer, construct a secondary set of layers with numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{2}{\varepsilon},\frac{2}{\varepsilon}.$
Note that this sequence is almost identical to that of the initializer except we keep going until a $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ is formed by the last two layers. The first layer in the secondary set is the same as the last layer of the initializer, each vertex of which is connected with $\frac{1}{\varepsilon}$ frozen white vertices and $\frac{1}{\varepsilon}$ frozen red vertices. Note the following observation:
Proposition 2.
No matter what happens in future layers, each of the $\frac{1}{\varepsilon}$ vertices in the last layer of the initializer can freely change colors.
Proof.
Each vertex in this layer has total degree $\frac{2}{\varepsilon}+1$ and will always have at least $\frac{1}{\varepsilon}$ frozen neighbors of every color, which is enough to play an uncertain best response no matter what the current color is. ∎
It is our careful construction of the initializer to get to this point combined with Proposition 2 that is the key observation needed to improve the bound in Balcan et al. (2011). We now finish the construction. Note that the $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ produced at the end of the secondary set is a starting point from which we can perform the boosting operations as defined above. Here we add a layer of $\frac{2}{\varepsilon}+1$ vertices to form an initial $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}+1}$ and then initialize $m$ to be $\frac{2}{\varepsilon}+1$. Starting from this layer, layers will increase in size by a constant multiplicative factor of $1+\frac{\varepsilon}{2}$ until we reach $c\cdot\varepsilon n$ vertices in the last layer for some small constant $c$ to be set later. The total number of vertices used in the boosting operations is $\Theta\left(\frac{c\cdot\varepsilon n}{1-\frac{1}{1+\frac{\varepsilon}{2}}}\right)=\Theta(cn)$. Combined with the number of vertices $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$ used in the initializer as well as the secondary set, we see that for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$ and a constant $c$ independent of $\varepsilon$ and $n$ that the total number of vertices used is at most $n$, which means our construction has the right number of vertices.
How many bad edges does our construction yield? Starting from the first layer in the secondary set, we can turn each successive layer red in sequence so that the last two layers form a $K_{c\cdot\varepsilon n,c\cdot\varepsilon n}$ where the left side is red and the right half is white. However, this produces only $c^{2}\cdot\varepsilon^{2}n^{2}$ bad edges, which, recalling that we started with $\frac{1}{\varepsilon}$ bad edges, matches the bound $\Omega(\varepsilon^{3}n^{2})$ on the price of uncertainty in Balcan et al. (2011). It turns out we can produce many more bad edges in our construction by doing the following. Note that the last layer in our construction is white, and all layers between the frozen layers defined above and the last layer are now red. We now turn each vertex in the layer after the frozen layers white, which is a valid uncertain best response by Proposition 2. We can now turn each successive layer white in sequence until the last three layers are white, red, and white respectively. Repeatedly switching the colors of the vertices in the layer after the frozen layers and propagating uncertain best responses, we can reach a state in which the boosting operation layers alternate between white and red. Every edge between boosting operation layers is now bad, which makes $\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{1-\left(\frac{1}{1+\frac{\varepsilon}{2}}\right)^{2}}\right)=\Theta(\varepsilon n^{2})$ bad edges at the ending state. Recalling that we began with $\frac{1}{\varepsilon}$ bad edges, we have established that the price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ in the domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
3 Upper Bound
We first provide a simple proof, omitted in Balcan et al. (2011), that the price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
Theorem 3 ((Balcan et al., 2011, Theorem 2)).
The price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
Proof.
We prove that in order to increase the number of bad edges, there must have been $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start. Consider the first step at which the number of bad edges increases. Note that in order for a vertex’s uncertain best response to increase the number of bad edges, the vertex must think it has more bad neighbors than it actually has, hence the number of bad edges must be $\Omega\left(\frac{1}{\varepsilon}\right)$. Since this is the first step at which the number of bad edges increases, the number of bad edges at the start must be $\Omega\left(\frac{1}{\varepsilon}\right)$. Since there can be at most $\binom{n}{2}=O(n^{2})$ bad edges at any step, we conclude that the price of uncertainty is $O(\varepsilon n^{2})$. ∎
3.1 Ideas for improving the upper bound
The second main result of this paper is that the true price of uncertainty is $O(\varepsilon^{2}n^{2})$.
Theorem 4.
The price of uncertainty for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$.
Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process. Consider a time in which there are $m^{2}$ bad edges. The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$. If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$. However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response. Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$. Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again. Assuming most edges can eventually be turned bad again as is the case in Section 2, the total number of bad edges at the end is
$O\left(\varepsilon^{2}n^{2}+\frac{\varepsilon^{2}n^{2}}{1+\varepsilon}+\cdots\right)=O(\varepsilon n^{2}).$
Since there must be $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start as in Theorem 3, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
Motivated by the intuition in Section 3.1, it appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses. The hope is that we can prove an upper bound on the growth rate in line with the rate described in Section 3.1. Unfortunately, this approach fails, as illustrated by the following example.
Example 5.
Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega\left(\frac{1}{\varepsilon}\right)$. The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$. We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
  • Let $V$ play an uncertain best response that switches its color to red, resulting in $m+1$ bad edges.
  • Take the endpoint of one such bad edge adjacent to $W$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $W$.
  • Switch the color of $V$ again, resulting in $m+1$ bad edges. Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $W$, and 1 bad edge incident to $R$.
  • Repeat the above, at each step switching the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once. There are now $m$ bad edges incident to $R$ and $m+1$ bad edges incident to $W$.
Note that Example 5 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges. The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in Section 3.1 it is $1+\varepsilon$. Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step. We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the time step at which each vertex first plays an uncertain best response. This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last. Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad. If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average. This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” as a clever monovariant.
3.2 Proof of Improved Upper Bound
We now give the formal proof of Theorem 4. It will be important to first establish a bound upon the starting number of bad edges.
Lemma 6.
In order to obtain a price of uncertainty on the order of $\Omega\left(\varepsilon^{2}n^{2}\right)$, the starting number of bad edges must be equal to $O\left(\varepsilon^{-2}\right)$.
Proof.
There can be at most $O\left(n^{2}\right)$ bad edges in a graph with $n$ vertices, so the starting number of bad edges must be
$\frac{O\left(n^{2}\right)}{\Omega\left(\varepsilon^{2}n^{2}\right)}=O\left(\varepsilon^{-2}\right).$
Let $G$ be the set of vertices of a graph upon which a consensus game is being played. Define $S_{0}$ as the set of vertices in $G$ that begin incident to at least one bad edge. Consider the set of vertices that begin without any bad edges and play at least one uncertain best response. They can be numbered in order of their first uncertain best response from first to last as $v_{1},v_{2},\dots,v_{m}$. Define a set $V$ as $\{v_{1},v_{2},\dots,v_{m}\}$. Finally, let $L$ be the set of vertices that begin without any bad edges and never play an uncertain best response. Notice that $S_{0}$, $V$, and $L$ are pairwise disjoint, and $S_{0}\cup V\cup L=G$.
For all integers $1\leq k\leq m$, let $S_{k}=S_{k-1}\cup\{v_{k}\}$.
Lemma 7.
Let $k$ be an integer between 0 and m1m-1. Before $v_{k+1}$ plays its first uncertain best response, every bad edge is incident to at least one vertex in $S_{k}$. Furthermore, every bad edge is always incident to at least one vertex in $S_{m}$.
Proof.
Before any uncertain best responses are played, all bad edges are incident to vertices in $S_{0}$ by definition. Similarly, at this point, edges between two vertices in $G\setminus S_{0}$ must all be good.
Let $0\leq j\leq m-1$ be an integer. Suppose that Lemma 7 holds when $k=j$. Then, any edge between two vertices in $G\setminus S_{j}$ is good. By construction, the first vertex outside of $S_{j}$ to play an uncertain best response is $v_{j+1}$. When this happens, the only edges to be affected will be those incident to $v_{j+1}$. However, since $v_{j+1}\in S_{j+1}$, every bad edge must now be incident to at least one vertex in $S_{j+1}$. Finally, as long as no vertices outside of $S_{j+1}$ change colors, every bad edge will remain incident to at least one vertex in $S_{j}$. By induction, Lemma 7 must hold for $k$ between 0 and m1m-1.
Finally, by construction, all vertices in $G\setminus S_{m}$ never play an uncertain best response. This means that any edges incident to 2 of these vertices will remain good, so Lemma 7 holds for $k=m$ as well. ∎
Lemma 8.
If $\varepsilon=\omega\left(n^{-1/2}\right)$, it is optimal to have $m=\Theta(n)$.
Proof.
Clearly, $m=|V|\leq|G|=n$, so $m=O(n)$, so the upper bound holds. We will now attack the lower bound.
By construction, $G$ is comprised of the two disjoint sets $S_{m}$ and $L$. By Lemma 7, any vertices in $L$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge. This means that we can delete these vertices along with all of their incident edges, which will always increase the final price of uncertainty. By construction, prior to when $v_{1}$ plays its first uncertain best response, any edges incident to 2 vertices in $G\setminus S_{0}$ must be good. Since $L=G\setminus S_{m}\subset G\setminus S_{0}$ and since vertices in $L$ never change their colors, any edges incident to two vertices in $L$ must always be good. By removing all such edges, we will always increase the price of uncertainty.
We are now left with a new set of vertices $L^{\prime}\subset L$ where every edge incident to a vertex in $L^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=V$. For $1\leq k\leq m$, $v_{k}$ can only play its first uncertain best response if its number of bad edges is at least its number of good edges multiplied by $(1+\varepsilon)^{-1}$. By Lemma 7, the number of bad edges incident to $v_{k}$ is at most $|S_{k-1}|\leq|S_{m}|$. Right before $v_{k}$ plays its first uncertain best response, all edges incident to $v_{k}$ and a vertex in $L^{\prime}$ must be good, so
$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|L^{\prime}|$
$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|L^{\prime}|$
$\Theta\left(|S_{m}|\right)$ $=\Theta\left(L^{\prime}\right).$
Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$. Recall that $|S_{m}|=|S_{0}|+m$. By Lemma 6, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin incident to a bad edge is at most twice this quantity. So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.
$\Theta(n)$ $=|S_{m}|$
$\Theta(n)$ $=|S_{0}|+m$
$\Theta(n)$ $=O\left(\varepsilon^{-2}\right)+m$
Since $\varepsilon=\omega\left(n^{-1/2}\right)$, this becomes
$\Theta(n)$ $=o\left(n\right)+m$
$\Theta(n)$ $=m.$
With our construction, Lemma 7 allows us to contain the bad edges in the graph inside one of $S_{0},S_{1},\dots,S_{m}$. We will now present a combinatorial abstraction that will allow us to focus on the effect of the first uncertain best response played by vertices in $V$.
We can use $S_{k}$ to construct an infinite sequence $E_{k}$ of positive integers, finitely many of which are nonzero. Each nonzero element of $E_{k}$ will encode the number of edges incident to a vertex in $G\setminus S_{k}$ and any vertex in $S_{k}$, and each vertex in $G\setminus S_{k}$ will be encoded exactly once in such a manner. The rest of $E_{k}$ will be populated with zeros. Finally, to ensure that $E_{k}$ is unique, we will also include a requirement that $E_{k}$ must be sorted in descending order. For example, suppose that there existed 4 edges incident to a vertex in $S_{0}$ and some vertex $x_{1}$, and 3 edges incident to a vertex in $S_{0}$ and some vertex $x_{2}\neq x_{1}$. Also, suppose that there existed no other edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$. Then, we would have
$E_{0}=\{4,3,0,0,0,\dots\}.$
Consider the state of the graph right before $v_{k}$ plays its first uncertain best response. By Lemma 7, all edges incident to $v_{k}$ and a vertex in $G\setminus S_{k-1}$ must be good. If we let $z$ be the number of edges incident to $v_{k}$ and a vertex in $S_{k-1}$, then the degree of $v_{k}$ cannot be greater than $z+\lfloor(1+\varepsilon)z\rfloor$. From this logic, going from $E_{k-1}$ to $E_{k}$ consists of deleting the element of $E_{k-1}$ corresponding to edges incident to a vertex in $S_{k-1}$ and $v_{k}$, which is $z$. We will now select $\lfloor(1+\varepsilon)z\rfloor$ elements of $E_{k-1}$ and increment them by $1$. This represents the new outgoing edges from $v_{k}$ to $G\setminus S_{k}$. Since any graph and sequence of uncertain best responses can be represented using the $E_{k}$ in such a manner, it suffices to study the behavior of these moves.
Definition 9.
A move on a sequence $E_{k}$ consists of the following steps:
  • Select some integer $z$ from $E_{k}$.
  • Delete one instance of $z$ from $E_{k}$.
  • Create $E_{k+1}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $E$ and adding 1 to each of them.
For the rest of this proof, let $\sum E_{k}$ denote the sum of the elements of $E_{k}$.
Lemma 10.
If $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$, then the final number of bad edges in the graph is $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ for all $\varepsilon=\omega\left(n^{-1/2}\right)$.
Proof.
Let $T_{k}$ denote the number of edges that are incident to two vertices in $S_{k}$. Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum E_{k}$ denote $(\sum E_{k})-(\sum E_{k-1})$. For every move that we play on an element $z$ in $E_{k}$, the sum $\sum E_{k}$ grows by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges that are contained inside of $S_{k}$ grows by $z$. This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$. We have
$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum E_{k}$
$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum E_{k}$
$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum E_{k}$
By summing both sides from $k=1$ to $m$, we get
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum E_{m}-\sum E_{0}.$
This lemma requires that $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$.
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}m^{2}|S_{0}|\right)-\sum E_{0}$
By Lemma 8, this becomes
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}+\Theta(n)+\varepsilon T_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)+\Theta(n)+\varepsilon T_{0}$
$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+\Theta\left(\varepsilon^{-1}n\right)+T_{0}.$
Since $\Omega\left(\varepsilon^{-1}\right)$ bad edges are required to play the first uncertain best response, we must have $S_{0}=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon n^{2}\right)$ must also be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. If $\varepsilon=\omega\left(n^{-1/2}\right)$, then $O\left(\varepsilon n^{2}\right)$ will dominate $\Theta\left(\varepsilon^{-1}n\right)$.
$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+T_{0}.$
Any vertex in $G$ is either in $S_{m}$, or $L$. This means that the only bad edges that are not counted by $T_{m}$ must be incident to a vertex in $S_{m}$ and a vertex in $L$. The edges that are incident to a vertex in $S_{m}$ and $G\setminus S_{m}=L$ are entirely described by $E_{m}$. Thus, the number of bad edges that exist between a vertex in $S_{m}$ and a vertex in $L$ is given by $\sum E_{m}$, which is $O\left(\varepsilon^{2}m^{2}\right)$. This is dominated by $T_{m}=O\left(\varepsilon n^{2}\right)+T_{0}$, so the asymptotic growth rate of the number of bad edges in the graph is given solely by that of $T_{m}$.
It remains to show that $T_{0}$ cannot grow faster than $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)\geq\Theta(\varepsilon n^{2})$. By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$. By Lemma 6, we start with $O\left(\varepsilon^{-2}\right)$ bad edges, so $T_{0}=O\left(\varepsilon^{-2}\right)$. This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega\left(n^{-1/2}\right)\geq\Theta\left(n^{-2/3}\right)$. ∎
Lemma 11.
The quantity $\sum E_{0}$ grows on the order of $O\left(m|S_{0}|\right)$.
Proof.
Since there are only $n$ vertices in the graph, every vertex in $S_{0}$ can have at most $n$ edges. Recall that $\sum E_{0}$ counts the number of edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$. By Lemma 8, we have
$\sum E_{0}$ $=O\left(n|S_{0}|\right)$
$\sum E_{0}$ $=O\left(m|S_{0}|\right).$
We will now consider the time reversed process of playing moves on the $E_{k}$.
Definition 12.
A reversed move on a sequence $E_{k}$ consists of the following steps:
  • Select a list of positive elements $\{e_{1},e_{2},\dots,e_{\alpha}\}$ from $E_{k}$
  • Subtract one from each of the $e_{j}$ for $1\leq j\leq\alpha$
  • Calculate $z=\lceil\alpha/(1+\epsilon)\rceil$
  • Create $E_{k-1}$ by inserting $z$ into $E_{k}$ as a new element
Let $\sum E_{k}^{2}$ denote the sum of the squares of the elements of $E_{k}$. In order to investigate the asymptotic growth rate of $\sum E_{m}$, we wish to find the largest starting value of $\sum E_{m}$ for which it is possible to play $m$ reversed moves to get to a desired value of $\sum E_{0}$.
Our plan is to bound the behavior of $\sum E_{k}$ and $\sum E_{k}^{2}$ as reversed moves are played. Intuitively, playing a reversed move on many elements of $E_{k}$ would result in a large decrease in from $\sum E_{k}$ to $\sum E_{k-1}$. However, playing such a move would also result in the creation of a new, very large element in $E_{k-1}$. Since we can only ever subtract one from each element per move, having the sum $\sum E_{k-1}$ concentrated in a few large elements will make decreasing $\sum E$ more difficult. We consider $\sum E_{k}^{2}$ as a way to measure how ”spread out” the sequence $E_{k}$ is, and this particular quantity is preferable over other measures because of how the change from $\sum E_{k}^{2}$ to $\sum E_{k-1}^{2}$ can be cleanly bounded by an expression in terms of $\sum E_{k}$.
Consider some arbitrary sequence $E_{k}$. Let $\alpha$ denote the number of elements of $E$ that we play the current reversed game move on, and let these elements be denoted $e_{1},e_{2},\dots,e_{\alpha}$. By the rules of the reversed game, playing such a move will create a new element of size $z=\lceil\alpha/(1+\varepsilon)\rceil$. It is easy to see that the effect of such a move on $\sum E_{k}$ is
$\sum E_{k-1}=\sum E_{k}-\alpha+z.$
Next, we must have
$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-\sum_{j=1}^{\alpha}\left(e_{j}^{2}-(e_{j}-1)^{2}\right)+z^{2}$
$=\sum E_{k}^{2}-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}.$
Since $e_{1},e_{2},\dots,e_{\alpha}$ are all elements of $E_{k}$, their sum must be bounded by $\sum E_{k}$. By plugging this into our expression for the change in $\sum E^{2}$, we get
$\sum E_{k-1}^{2}-\sum E_{k}^{2}$
$={}$ $-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}$
$\geq{}$ $-2\sum_{j=1}^{\alpha}e_{j}+z^{2}$
$\geq{}$ $-2\sum E_{k}+\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil^{2}$
$\geq{}$ $-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}$ (1)
Lemma 13.
It is always optimal to take the lower bound in (1) for the change in $\sum E^{2}$.
Proof.
Notice that neither our change in $\sum E$ nor our bound (1) on the change in $\sum E^{2}$ depend on $\sum E_{k}^{2}$. If we make $\sum E_{k}^{2}$ as small as possible for a fixed $\sum E_{k}$, the future options we will have for $\alpha$ will not change. Furthermore, we only care that $\sum E_{0}^{2}$ ends up being smaller than $\left(\sum E_{0}\right)^{2}$. This means that allowing $\sum E_{k-1}^{2}$ to attain its lower bound will never result in more moves being required. There are no further issues because there is no need that intermediate values can be represented by actual sequences; there is no required minimum value for $\sum E_{k}^{2}$ for any integer $0\leq k\leq m-1$. ∎
We can also obtain a simpler lower bound for the change in $\sum E$.
$\sum E_{k-1}-\sum E_{k}$
$={}$ $-\alpha+z$
$={}$ $\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
$\geq{}$ $(1-\varepsilon)\alpha-\alpha$
$={}$ $-\varepsilon\alpha.$ (2)
Lemma 14.
It is always optimal to take the lower bound in (2) for the change in $\sum E$.
Proof.
If we allow $\alpha\in\mathbb{R}$, still with $\alpha\geq 1$, it is always possible to choose a new $\alpha^{\prime}$ such that
$\varepsilon\alpha^{\prime}=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha.$
The RHS of the previous equation is clearly nonpositive, so $\alpha^{\prime}$ must be nonpositive as well. We can also bound the size of $\alpha^{\prime}$ with
$-\varepsilon\alpha^{\prime}$ $=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
$1-\varepsilon\alpha^{\prime}$ $\geq 1+\frac{\alpha}{1+\varepsilon}-\alpha$
$-\varepsilon\alpha^{\prime}$ $\geq\frac{-\varepsilon\alpha}{1+\varepsilon}$
$(1+\varepsilon)\alpha^{\prime}$ $\leq\alpha$
$\alpha^{\prime}$ $<\alpha.$
By (1), if we take the lower bound for the change in $\sum E_{k}^{2}$ using $\alpha^{\prime}$, the value of $\sum E_{k-1}^{2}$ is equal to or smaller than if we had used $\alpha$ instead. By the same logic that was used to prove Lemma 13, it is always helpful to have a lower $\sum E^{2}$. ∎
With Lemma 13 and Lemma 14, we now have a strictly stronger set of actions on $\sum E$ and $\sum E^{2}$ that we can perform with each move.
Definition 15.
A strong reversed move that uses $\alpha$ on the pair of values $\sum E_{k}$ and $\sum E_{k}^{2}$ is given by
$\sum E_{k-1}$ $=\sum E_{k}-\varepsilon\alpha$
$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}.$
For integers $1\leq k\leq m$, let $\alpha_{k}$ denote the value of $\alpha$ used in the strong reversed move to get from $\left(\sum E_{k},\sum E_{k}^{2}\right)$ to $\left(\sum E_{k-1},\sum E_{k-1}^{2}\right)$. Then,
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\alpha_{1}+\varepsilon\alpha_{2}+\cdots+\varepsilon\alpha_{m}$
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}.$ (3)
The same process is a little more complicated for $\sum E_{m}^{2}$.
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
We can rewrite the $\sum E_{j}$ by plugging in (3).
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{0}+\varepsilon\sum_{i=1}^{j}\alpha_{i}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2m\sum E_{0}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $=\sum E_{0}^{2}+2m\sum E_{0}-\sum E_{m}^{2}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
We have $\sum E_{m}^{2}\geq\sum E_{m}$. Furthermore, due to (3), $\sum E_{m}\geq\varepsilon\sum_{j=1}^{m}\alpha_{j}$. We can use these two inequalities to replace $\sum E_{m}^{2}$.
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$ (4)
By the Cauchy-Schwarz Inequality, we have
$\left(\sum_{j=1}^{m}1\right)\left(\sum_{j=1}^{m}\alpha_{j}^{2}\right)$ $\geq\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
$\sum_{j=1}^{m}\alpha_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}.$
We can then plug this into (4):
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
$0$ $\geq(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}-\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-\left(\sum E_{0}^{2}+2m\sum E_{0}\right).$
We are now left with a quadratic in the sum $\sum_{j=1}^{m}\alpha_{j}$. Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}\alpha_{j}$ must be less than the larger root of this quadratic. This larger root is obtained through a tedious application of the quadratic formula:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\varepsilon(2m-1)+\sqrt{\varepsilon^{2}(2m-1)^{2}+4(1-\varepsilon)^{2}m^{-1}\left(\sum E_{0}^{2}+2m\sum E_{0}\right)}}{(1-\varepsilon)^{2}m^{-1}}.$ (5)
Lemma 16.
If $\varepsilon=\Omega\left(n^{-1/4}\right)$, then $\sum_{j=1}^{m}\alpha_{j}$ is bounded by $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
Proof.
If we substitute asymptotic behavior into the RHS of (5), we get
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum E_{0}^{2}\right)+\Theta\left(\sum E_{0}\right)}}{\Theta\left(m^{-1}\right)}.$
Notice that we must have $\sum E_{0}^{2}\leq\left(\sum E_{0}\right)$. By Lemma 11, we can rewrite the previous inequality as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)+\Theta\left(m|S_{0}|\right)}}{\Theta\left(m^{-1}\right)}$
It is clear that the $\Theta\left(m|S_{0}|^{2}\right)$ term dominates the $\Theta\left(m|S_{0}|\right)$ term. We now have the following:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}.$ (6)
Depending on which one of the two terms inside of the square root dominate, there are two cases that we have to tackle.
Case 16.1.
$\Theta\left(\varepsilon^{2}m^{2}\right)$ dominates $\Theta\left(m|S_{0}|^{2}\right)$.
In this case, we can rewrite (6) as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon m^{2}\right).$
Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
Case 16.2.
$\Theta\left(m|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}m^{2}\right)$.
In this case, we can rewrite (6) as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(m^{3/2}|S_{0}|\right).$
By Lemma 8, $\varepsilon=\Omega\left(n^{-1/4}\right)$ is equivalent to $\varepsilon=\Omega\left(m^{-1/4}\right)$.
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}m^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(m^{-1/4}\right)^{-2}m^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$
The result of Lemma 16 can be plugged into (3) with Lemma 11 to obtain a bound on $\sum E_{m}$. Since we are only interested in $\varepsilon=\Omega\left(n^{-1/4}\right)\geq\Theta\left(n^{-1/3}\right)$, we have
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}$
$\sum E_{m}$ $=O\left(m|S_{0}|\right)+O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$
$\sum E_{m}$ $=O\left(\varepsilon^{3}m^{2}|S_{0}|\right).$ (7)
By (7) and Lemma 10, the final number of bad edges in the graph must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. By construction, every vertex in $S_{0}$ begins with at least one bad edge, so we must begin with $\Omega\left(|S_{0}|\right)$ bad edges. We obtain the following upper bound on the price of uncertainty for consensus games when $\varepsilon=\Omega\left(n^{-1/4}\right)$:
$\frac{O\left(\varepsilon^{2}n^{2}|S_{0}|\right)}{\Omega\left(|S_{0}|\right)}=O\left(\varepsilon^{2}n^{2}\right).$
This finishes the proof of Theorem 4.

### v2

@@
-Price of Uncertainty for Consensus Games
+The Price of Uncertainty for Social Consensus
 Abstract
-Many game-theoretic models assume that players have access to accurate information, but uncertainty in observed data is frequently present in real-world settings.
-In this paper, we consider a model of uncertainty where adversarial perturbations of relative magnitude $1+\varepsilon$ are introduced to players’ observed costs.
-The effect of uncertainty on social cost is referred to as the price of uncertainty.
-We prove a tight bound on the price of uncertainty for consensus games of $\Theta(\varepsilon^{2}n^{2})$ for all $\varepsilon=\Omega\mathopen{}\left(n^{-1/4}\right)$.
-This improves a previous lower bound of $\Omega(\varepsilon^{3}n^{2})$ as well as a previous upper bound of $O(\varepsilon n^{2})$.
+How hard is it to achieve consensus in a social network under uncertainty?
+In this paper we model this problem as a social graph of agents where each vertex is initially colored red or blue.
+The goal of the agents is to achieve consensus, which is when the colors of all agents align.
+Agents attempt to do this locally through steps in which an agent changes their color to the color of the majority of their neighbors.
+In real life, agents may not know exactly how many of their neighbors are red or blue, which introduces uncertainty into this process.
+Modeling uncertainty as perturbations of relative magnitude $1+\varepsilon$ to these color neighbor counts, we show that even small values of $\varepsilon$ greatly hinder the ability to achieve consensus in a social network.
+We prove theoretically tight upper and lower bounds on the price of uncertainty, a metric defined by Balcan et al.
+(2009) to quantify the effect of uncertainty in network games.
 1 Introduction
-If you are asked whether more of your friends like vanilla ice cream or chocolate ice cream, could you truly give the right answer with confidence?
-Even the best polling organizations sometimes fail to predict the outcomes of two-party majority elections.
-The truth is that uncertainty is everywhere and can impact our abilities to make rational decisions.
-However, most game-theoretic models assume that players have access to accurate information about interactions as well as their own costs.
-In the context of strategic interactions, it is interesting to see how uncertainty can affect agents’ behavior.
+In the past decade the rapid expansion of social media platforms has resulted in the formation of large and highly connected social networks.
+The evolution of these networks has been extensively studied in order to better understand how opinions held by agents are affected by their interactions with other people.
+In a social setting, individuals generally tend to modify their beliefs to more closely align with others that they frequently interact with (Moussaïd et al., 2013; Garimella et al., 2018).
+This generally causes opinions held by agents in a network to generally converge over time as interacting agents attempt to reach a consensus.
+However, previous research has shown that additional factors, such as a sparse network structure or agents who refuse to change their opinion, can hinder consensus formation.
+(Acemoğlu et al., 2013; Hegselmann and Krause, 2002; Moussaïd et al., 2013).
+The vast majority of social network models in the existing literature assume that agents have accurate information about the opinions held by others.
+However, in large networks where vertices generally have many neighbors and where interactions between agents are often brief, agents are often unable to make accurate judgments about the opinions of other agents such as the number of their consenting or dissenting peers (Xiong and Liu, 2014).
+Previous work modeling uncertainty in games has focused on the failure of agents to always pick the action with the optimal payoff.
+For example, in the quantal response model, agents sometimes deviate from optimal behavior with strategies providing higher payoffs being chosen more often than those providing lower payoffs (Penna, 2018; Blume, 1993; Alós-Ferrer and Netzer, 2010).
+Alternatively, agents may occasionally fail to participate in the game, which can sometimes increase the social cost (Meir et al., 2012; Li et al., 2017).
+In this paper, we instead focus on the effect of uncertainty in agents’ observations rather than their actions.
+We model a social network as a graph with an agent at each vertex.
+Individuals have a choice between two colors, red or blue, and the goal of each agent is to be the same color as the majority of their neighbors.
+This is called a consensus game (Christodoulou et al., 2012; Fabrikant et al., 2004).
+We consider a model of uncertainty first proposed by Balcan et al.
+(2009) where agents experience small, adversarial perturbations of magnitude $1+\varepsilon$ to their observations of their neighbors.
+These perturbations capture the difficulty of ascertaining accurate information in large social networks, as the absolute margin of error increases proportionally to the number of neighbors an agent has.
+Alternatively, this model can also capture the effect of biased measurements, where agents may distort observations to agree with their beliefs (confirmation bias) or place undue emphasis on new information (novelty bias) (Ngampruetikorn and Stephens, 2016; Mao et al., 2024).
+The goal of this paper is to study how uncertainty affects behavior in consensus games as agents make potentially suboptimal updates to their colors.
 Balcan et al.
-(2009) consider how introducing small perturbations in observed costs in well-studied potential games can impact their behavior under improved-response and best-response dynamics.
-They coin the term price of uncertainty, and study it in a wide variety of potential games including set-covering games (Buchbinder et al., 2008), fair cost-sharing games (Anshelevich et al., 2008), job-scheduling games (Nisan et al., 2007), and consensus games.
-In some games, even extremely small perturbations can cause these dynamics to snowball out of control and move to states of much higher social cost, whereas in other games these dynamics are much more stable against large degrees of perturbation.
-Consensus games are a special case of party affiliation games, which are in turn a class of potential games (Bhalgat et al., 2010).
-In a consensus game, each of $n$ players can choose between two parties: the white group and the red group.
-Players accrue a constant positive cost for interacting with those that chose the other side.
-Players are modeled as vertices of an undirected graph that can be colored either white or red, and a player’s cost is the number of neighbors with the other color, which we call bad neighbors.
+(2009) define a social cost function as the number of edges whose endpoints have different colors.
+They quantify the effect of uncertainty using a measure called the price of uncertainty (PoU) of consensus games, defined as the maximum multiplicative factor by which the social cost can increase from any starting state.
+Even extremely small perturbations can allow adversarial perturbations to snowball out of control and cause the game to move to states of much higher social cost (Balcan et al., 2011).
+Our main results are tight asymptotic upper and lower bounds on the PoU for consensus games.
+Letting $n$ denote the number of agents participating in the consensus game, the previously best known asymptotic bounds due to Balcan et al.
+(2011) were a lower bound of $\Omega(n^{2}\varepsilon^{3})$ and an upper bound of $O(n^{2}\varepsilon)$.
+In this paper, we extend their work by proving that the PoU for consensus games is exactly $\Theta(n^{2}\varepsilon^{2})$.
+We believe that our result also holds irrespective of agent update ordering (i.e., the adversary can only control the magnitude of the perturbations), and this a potential future direction for research.
+2 The Model
+We consider a consensus game, where each of $n$ agents are located at distinct vertices of a simple undirected graph with $n$ vertices (Christodoulou et al., 2012).
+In this paper, we will use the words agent and vertex interchangeably.
+Each agent can choose a color from $\{0,1\}$.
+Let us refer to two agents as neighbors iff there exists an edge between the two agents.
+An agent’s cost is defined as the number of their neighbors that pick a different color, which we call bad neighbors.
 We will also call edges whose endpoints have different colors bad edges.
-Starting from a ground state with social cost at least 1, the price of uncertainty is the maximum multiplicative factor by which the social cost can increase via uncertain best-response dynamics.
-The term uncertain here refers to the fact that a player can miscount their cost, which recall is the number of bad neighbors, by a multiplicative factor of at most $1+\varepsilon$ for a small uncertainty factor $\varepsilon>0$ that can depend on $n$.
-Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise a player cannot miscount its number of bad neighbors and hence there is no uncertainty.
-Balcan et al.
-(2011) prove an upper bound of $O(\varepsilon n^{2})$ on the price of uncertainty for consensus games that holds for all $\varepsilon=\Omega\left(\frac{1}{n}\right)$ as well as a lower bound of $\Omega(\varepsilon^{3}n^{2})$ that holds in the domain $\varepsilon=\Omega\left(n^{-1/3}\right)$.
-This paper obtains a lower bound for the price of uncertainty of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
-This is achieved using a construction that amplifies the snowball effect of uncertainty described by Balcan et al.
+Finally, for any state $S$ of the game, we define a social cost function $\operatorname{cost}(S)$ as the sum of the costs of every agent in the game.
+Starting from a ground state $S_{0}$ with social cost at least 1, let $S_{t}$ be a state achieved after $t$ steps under uncertain best-response dynamics.
+Here, the term uncertain here refers to the fact that an agent might miscount their cost by a multiplicative factor of at most $1+\varepsilon$ for an uncertainty factor $\varepsilon>0$ that can depend on $n$.
+Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty.
+By definition, an uncertain best response must involve a color switch.
+Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise an agent cannot miscount its number of bad neighbors and hence there is no uncertainty.
+The price of uncertainty (PoU) of consensus games is defined as
+$\operatorname{PoU}(\varepsilon,\text{consensus})=\max\left(\frac{\operatorname{cost}(S_{t})}{\operatorname{cost}(S_{0})}\right),$
+where the maximum is taken over all initial states $S_{0}$ and all states $S_{t}$ reachable from $S_{0}$ in some number $t$ of steps (Balcan et al., 2009).
+This paper obtains a lower bound for the PoU of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
+Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
+This is achieved using an intricate construction that amplifies the “snowball” effect of uncertainty in Balcan et al.
 (2011).
-Then, we will prove that the price of uncertainty for consensus games is also bounded by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega\left(\varepsilon^{-1/4}\right)$ through the construction and analysis of a clever monovariant.
-These bounds coincide, so we conclude that the price of uncertainty for consensus games is $\Theta\left(\varepsilon^{2}n^{2}\right)$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$, resolving an open problem proposed by Balcan et al.
+Next, we prove that the PoU for consensus games is also bounded from above by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega(\varepsilon^{-1/4})$ by analyzing a clever monovariant.
+These bounds coincide, so we conclude that the asymptotically tight PoU for consensus games is $\Theta(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$, resolving an open problem of Balcan et al.
 (2011).
+2.1 Main Technical Challenges
+Lower Bound.
+In order to achieve a lower bound of $\Omega(\varepsilon^{3}n^{2})$, Balcan et al.
+(2011) utilizes an initializer gadget to set up initial colorings for an output component.
+This component consists of successive layers of vertices, where each layer has more vertices than the previous layer by a factor of $(1+\varepsilon)$, and every vertex is connected to all vertices on adjacent layers.
+This output component is responsible for generating $\Omega(\varepsilon n^{2})$ bad edges, and the initializer gadget requires $\Theta(1/\varepsilon^{2})$ starting bad edges, so this achieves a lower bound of $\Omega(\varepsilon^{3}n^{2})$.
+We have found that it is very difficult to increase the final number of bad edges produced by such an output for any construction because the graph immediately becomes too connected for enough agents to play uncertain best responses.
+Rather, we present an alternative construction for an initializer gadget in Section 3.1 that only requires $\Theta(1/\varepsilon)$ starting bad edges, which results in a construction for which the number of bad edges can increase by a factor of $O(\varepsilon^{2}n^{2})$.
+Upper Bound.
+Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process.
+Consider a time in which there are $m^{2}$ bad edges.
+The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$.
+If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$.
+However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response.
+Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
+For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$.
+Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again.
+Assuming most edges can eventually be turned bad again as is the case in our lower bound construction, the total number of bad edges at the end is
+$O\left(\varepsilon^{2}n^{2}+\frac{\varepsilon^{2}n^{2}}{1+\varepsilon}+\cdots\right)=O(\varepsilon n^{2}).$
+Since there must be $\Omega(1/\varepsilon)$ bad edges at the start, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
+It appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses.
+The hope is that we can prove an upper bound on the growth rate in line with our estimate of $O(\varepsilon^{2}n^{2})$.
+Unfortunately, this approach fails, as illustrated by the following example.
+Example 1.
+Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega(1/\varepsilon)$.
+The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$.
+We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
+  • Let $V$ play an uncertain best response that switches its color to red, resulting in $m+1$ bad edges.
+  • Take the endpoint of one such bad edge adjacent to $W$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $W$.
+  • Switch the color of $V$ again, resulting in $m+1$ bad edges.
+Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $W$, and 1 bad edge incident to $R$.
+  • Repeat the above, at each step switching the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once.
+There are now $m$ bad edges incident to $R$ and $m+1$ bad edges incident to $W$.
+Note that Example 1 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges.
+The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in our initial estimate it was $1+\varepsilon$.
+Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step.
+We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
+Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the step at which each vertex first plays an uncertain best response.
+This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last.
+Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad.
+If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average.
+This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
+Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” as a clever monovariant.
 3 Lower Bound
-We first improve the lower bound for the price of uncertainty for consensus games.
-Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
+We first improve the lower bound for the PoU for consensus games.
+Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
 Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
 Theorem 2.
-The price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
+The PoU for consensus games is $\Omega(\varepsilon^{2}n^{2})$ for $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
 To prove their lower bound, Balcan et al.
@@
 3.1 Proof of Improved Lower Bound
-Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty.
-By definition, an uncertain best response must involve a color switch.
 One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level.
@@
 4 Upper Bound
-We first provide a simple proof, omitted in Balcan et al. (2011), that the price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
-Theorem 3 ((Balcan et al., 2011, Theorem 2)).
-The price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
-Proof.
-We prove that in order to increase the number of bad edges, there must have been $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start.
-Consider the first step at which the number of bad edges increases.
-Note that in order for a vertex’s uncertain best response to increase the number of bad edges, the vertex must think it has more bad neighbors than it actually has, hence the number of bad edges must be $\Omega\left(\frac{1}{\varepsilon}\right)$.
-Since this is the first step at which the number of bad edges increases, the number of bad edges at the start must be $\Omega\left(\frac{1}{\varepsilon}\right)$.
-Since there can be at most $\binom{n}{2}=O(n^{2})$ bad edges at any step, we conclude that the price of uncertainty is $O(\varepsilon n^{2})$.
-∎
-3.1 Ideas for improving the upper bound
 The second main result of this paper is that the true price of uncertainty is $O(\varepsilon^{2}n^{2})$.
 Theorem 4.
-The price of uncertainty for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$.
-Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process.
-Consider a time in which there are $m^{2}$ bad edges.
-The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$.
-If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$.
-However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response.
-Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
-For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$.
-Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again.
-Assuming most edges can eventually be turned bad again as is the case in Section 2, the total number of bad edges at the end is
-$O\left(\varepsilon^{2}n^{2}+\frac{\varepsilon^{2}n^{2}}{1+\varepsilon}+\cdots\right)=O(\varepsilon n^{2}).$
-Since there must be $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start as in Theorem 3, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
-Motivated by the intuition in Section 3.1, it appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses.
-The hope is that we can prove an upper bound on the growth rate in line with the rate described in Section 3.1.
-Unfortunately, this approach fails, as illustrated by the following example.
-Example 5.
-Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega\left(\frac{1}{\varepsilon}\right)$.
-The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$.
-We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
-  • Let $V$ play an uncertain best response that switches its color to red, resulting in $m+1$ bad edges.
-  • Take the endpoint of one such bad edge adjacent to $W$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $W$.
-  • Switch the color of $V$ again, resulting in $m+1$ bad edges.
-Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $W$, and 1 bad edge incident to $R$.
-  • Repeat the above, at each step switching the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once.
-There are now $m$ bad edges incident to $R$ and $m+1$ bad edges incident to $W$.
-Note that Example 5 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges.
-The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in Section 3.1 it is $1+\varepsilon$.
-Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step.
-We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
-Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the time step at which each vertex first plays an uncertain best response.
-This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last.
-Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad.
-If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average.
-This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
-Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” as a clever monovariant.
+The PoU for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$.
+In addition to focusing only on the steps at which each vertex plays its first uncertain best response, we will also carefully construct and analyze a monovariant that restricts the degrees of new vertices and allows us to bound the number of edges in the graph.
 4.1 Proof of Improved Upper Bound
-We now give the formal proof of Theorem 4. It will be important to first establish a bound upon the starting number of bad edges.
-Lemma 6.
-In order to obtain a price of uncertainty on the order of $\Omega\left(\varepsilon^{2}n^{2}\right)$, the starting number of bad edges must be equal to $O\left(\varepsilon^{-2}\right)$.
+Let $G$ be a graph for a consensus game, and let $V$ be the set of vertices of $G$. It will be useful to first establish a bound upon the starting number of bad edges.
+Proposition 5.
+Suppose that it is possible to increase the number of bad edges in $G$ by a factor of $\Omega\left(\varepsilon^{2}n^{2}\right)$.
+Then, the starting number of bad edges must be $O\left(\varepsilon^{-2}\right)$.
 Proof.
@@
 $\frac{O\left(n^{2}\right)}{\Omega\left(\varepsilon^{2}n^{2}\right)}=O\left(\varepsilon^{-2}\right).$
-Let $G$ be the set of vertices of a graph upon which a consensus game is being played.
-Define $S_{0}$ as the set of vertices in $G$ that begin incident to at least one bad edge.
-Consider the set of vertices that begin without any bad edges and play at least one uncertain best response.
-They can be numbered in order of their first uncertain best response from first to last as $v_{1},v_{2},\dots,v_{m}$.
-Define a set $V$ as $\{v_{1},v_{2},\dots,v_{m}\}$.
-Finally, let $L$ be the set of vertices that begin without any bad edges and never play an uncertain best response.
-Notice that $S_{0}$, $V$, and $L$ are pairwise disjoint, and $S_{0}\cup V\cup L=G$.
-For all integers $1\leq k\leq m$, let $S_{k}=S_{k-1}\cup\{v_{k}\}$.
-Lemma 7.
-Let $k$ be an integer between 0 and m1m-1.
-Before $v_{k+1}$ plays its first uncertain best response, every bad edge is incident to at least one vertex in $S_{k}$.
+We begin by partitioning $V$ into three disjoint subsets:
+  • Let $P$ be the set of vertices in $V$ that begin without any bad edges and play at least one best response.
+  • Let $Q$ be the set of vertices in $V$ that begin without any bad edges and never play a best response.
+  • Let $S_{0}$ be the set of all remaining vertices in $V$, i.e. those that begin incident to at least one bad edge.
+Let us number the vertices in $P$ in order of their first best responses from first to last as $p_{1},p_{2},\dots,p_{m}$ for $m=|P|$.
+Then, for all integers $1\leq k\leq m$, define $S_{k}=S_{0}\cup\{p_{1},p_{2},\dots,p_{k}\}$.
+Now that we have set up our notation, we will prove some preliminary statements regarding the behavior of the consensus game.
+Proposition 6.
+Let $k$ be an integer such that $0\leq k\leq m$.
+Before $p_{k+1}$ plays its first best response, every bad edge is incident to at least one vertex in $S_{k}$.
 Furthermore, every bad edge is always incident to at least one vertex in $S_{m}$.
 Proof.
-Before any uncertain best responses are played, all bad edges are incident to vertices in $S_{0}$ by definition.
-Similarly, at this point, edges between two vertices in $G\setminus S_{0}$ must all be good.
+Before any best responses are played, all bad edges have endpoints in $S_{0}$ by definition. Similarly, at this point, edges between two vertices in $V\setminus S_{0}$ must all be good.
 Let $0\leq j\leq m-1$ be an integer.
-Suppose that Lemma 7 holds when $k=j$.
-Then, any edge between two vertices in $G\setminus S_{j}$ is good.
-By construction, the first vertex outside of $S_{j}$ to play an uncertain best response is $v_{j+1}$.
-When this happens, the only edges to be affected will be those incident to $v_{j+1}$.
-However, since $v_{j+1}\in S_{j+1}$, every bad edge must now be incident to at least one vertex in $S_{j+1}$.
+Suppose that Proposition 6 holds when $k=j$.
+Then, any edge between two vertices in $V\setminus S_{j}$ is good.
+By construction, the first vertex outside of $S_{j}$ to play a best response is $p_{j+1}$.
+When this happens, the only edges to be affected will be those incident to $p_{j+1}$.
+However, since $p_{j+1}\in S_{j+1}$, every bad edge must now be incident to at least one vertex in $S_{j+1}$.
 Finally, as long as no vertices outside of $S_{j+1}$ change colors, every bad edge will remain incident to at least one vertex in $S_{j}$.
-By induction, Lemma 7 must hold for $k$ between 0 and m1m-1.
-Finally, by construction, all vertices in $G\setminus S_{m}$ never play an uncertain best response.
-This means that any edges incident to 2 of these vertices will remain good, so Lemma 7 holds for $k=m$ as well.
+By strong induction, Proposition 6 must hold for all integers $k$ between 0 and m1m-1 inclusive.
+Finally, by construction, all vertices in $V\setminus S_{m}=Q$ never play a best response.
+This means that any edges between 2 of these vertices will remain good, so Proposition 6 holds for $k=m$ as well.
 ∎
 Lemma 7.
-If $\varepsilon=\omega\left(n^{-1/2}\right)$, it is optimal to have $m=\Theta(n)$.
+Suppose that $\varepsilon=\omega(n^{-1/2})$. Then, the factor by which the number of bad edges in $G$ can be increased is maximized when $m=\Theta(n)$.
 Proof.
-Clearly, $m=|V|\leq|G|=n$, so $m=O(n)$, so the upper bound holds. We will now attack the lower bound.
-By construction, $G$ is comprised of the two disjoint sets $S_{m}$ and $L$.
-By Lemma 7, any vertices in $L$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge.
-This means that we can delete these vertices along with all of their incident edges, which will always increase the final price of uncertainty.
-By construction, prior to when $v_{1}$ plays its first uncertain best response, any edges incident to 2 vertices in $G\setminus S_{0}$ must be good.
-Since $L=G\setminus S_{m}\subset G\setminus S_{0}$ and since vertices in $L$ never change their colors, any edges incident to two vertices in $L$ must always be good.
-By removing all such edges, we will always increase the price of uncertainty.
-We are now left with a new set of vertices $L^{\prime}\subset L$ where every edge incident to a vertex in $L^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=V$.
-For $1\leq k\leq m$, $v_{k}$ can only play its first uncertain best response if its number of bad edges is at least its number of good edges multiplied by $(1+\varepsilon)^{-1}$.
-By Lemma 7, the number of bad edges incident to $v_{k}$ is at most $|S_{k-1}|\leq|S_{m}|$.
-Right before $v_{k}$ plays its first uncertain best response, all edges incident to $v_{k}$ and a vertex in $L^{\prime}$ must be good, so
-$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|L^{\prime}|$
-$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|L^{\prime}|$
-$\Theta\left(|S_{m}|\right)$ $=\Theta\left(L^{\prime}\right).$
+Clearly, $m=|P|<|V|=n$, so $m=O(n)$ and the upper bound holds. We will now attack the lower bound.
+By construction, $V$ consists of the two disjoint subsets $S_{m}$ and $Q$.
+By Proposition 6, any vertices in $Q$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge.
+This means that we can delete these vertices along with all of their incident edges, which will never decrease the factor by which the number of bad edges in $G$ can be increased.
+By construction, prior to when $p_{1}$ plays its first best response, any edges between 2 vertices in $V\setminus S_{0}$ must be good.
+Since $Q=V\setminus S_{m}\subset V\setminus S_{0}$, and since vertices in $Q$ never change their colors, any edges between two vertices in $Q$ must always be good.
+By removing all such edges, we will never decrease the factor by which the number of bad edges in $G$ can be increased.
+We are now left with a new set of vertices $Q^{\prime}\subset Q$ where every edge incident to a vertex in $Q^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=P$.
+For integers $1\leq k\leq m$, $p_{k}$ can only play its first best response if the number of its bad neighbors is at least the number of its good neighbors multiplied by $(1+\varepsilon)^{-1}$.
+By Proposition 6, the number of bad edges incident to $p_{k}$ is at most $|S_{k-1}|<|S_{m}|$.
+Right before $p_{k}$ plays its first best response, all edges between $p_{k}$ and a vertex in $Q^{\prime}$ must be good, so
+$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|Q^{\prime}|$
+$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|Q^{\prime}|$
+$\Theta\left(|S_{m}|\right)$ $=\Theta\left(Q^{\prime}\right).$
 Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$.
 Recall that $|S_{m}|=|S_{0}|+m$.
-By Lemma 6, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin incident to a bad edge is at most twice this quantity.
+By Proposition 5, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin with a bad neighbor is at most twice this quantity.
 So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.
@@
 $\Theta(n)$ $=m.$
-With our construction, Lemma 7 allows us to contain the bad edges in the graph inside one of $S_{0},S_{1},\dots,S_{m}$.
-We will now present a combinatorial abstraction that will allow us to focus on the effect of the first uncertain best response played by vertices in $V$.
-We can use $S_{k}$ to construct an infinite sequence $E_{k}$ of positive integers, finitely many of which are nonzero.
-Each nonzero element of $E_{k}$ will encode the number of edges incident to a vertex in $G\setminus S_{k}$ and any vertex in $S_{k}$, and each vertex in $G\setminus S_{k}$ will be encoded exactly once in such a manner.
-The rest of $E_{k}$ will be populated with zeros.
-Finally, to ensure that $E_{k}$ is unique, we will also include a requirement that $E_{k}$ must be sorted in descending order.
-For example, suppose that there existed 4 edges incident to a vertex in $S_{0}$ and some vertex $x_{1}$, and 3 edges incident to a vertex in $S_{0}$ and some vertex $x_{2}\neq x_{1}$.
-Also, suppose that there existed no other edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$.
+With our construction, Proposition 6 allows us to contain the bad edges in the graph at any step inside one of $S_{0},S_{1},\dots,S_{m}$.
+We will now present a combinatorial abstraction that will allow us to focus on the effect of the first best response played by each vertex in $P$.
+For every integer $k$ satisfying $0\leq k\leq m$, we can use $S_{k}$ to construct a multiset $A_{k}$ to encode information about certain bad edges in $S_{k}$.
+Specifically, each element of $A_{k}$ will describe the number of edges between a particular vertex in $V\setminus S_{k}$ and any vertex in $S_{k}$, and each vertex in $V\setminus S_{k}$ with at least one connection to a vertex in $S_{k}$ will be encoded exactly once in such a manner.
+Example 8.
+Suppose that there existed 4 edges between a vertex in $S_{0}$ and some vertex $x_{1}\not\in S_{0}$, and 3 edges between a vertex in $S_{0}$ and some different vertex $x_{2}\not\in S_{0}$.
+Also, suppose that there existed no other edges between a vertex in $S_{0}$ and a vertex in $V\setminus S_{0}$.
 Then, we would have
-$E_{0}=\{4,3,0,0,0,\dots\}.$
-Consider the state of the graph right before $v_{k}$ plays its first uncertain best response.
-By Lemma 7, all edges incident to $v_{k}$ and a vertex in $G\setminus S_{k-1}$ must be good.
-If we let $z$ be the number of edges incident to $v_{k}$ and a vertex in $S_{k-1}$, then the degree of $v_{k}$ cannot be greater than $z+\lfloor(1+\varepsilon)z\rfloor$.
-From this logic, going from $E_{k-1}$ to $E_{k}$ consists of deleting the element of $E_{k-1}$ corresponding to edges incident to a vertex in $S_{k-1}$ and $v_{k}$, which is $z$.
-We will now select $\lfloor(1+\varepsilon)z\rfloor$ elements of $E_{k-1}$ and increment them by $1$.
-This represents the new outgoing edges from $v_{k}$ to $G\setminus S_{k}$.
-Since any graph and sequence of uncertain best responses can be represented using the $E_{k}$ in such a manner, it suffices to study the behavior of these moves.
+$A_{0}=\{4,3\}.$
+$S_{0}$$x_{1}$$x_{2}$
+Figure 1: Diagram for Example 8.
+Consider the state of the graph right before $p_{k}$ plays its first best response.
+By Proposition 6, all edges between $p_{k}$ and a vertex in $V\setminus S_{k-1}$ must be good.
+If we let $z$ be the number of edges incident to $p_{k}$ and a vertex in $S_{k-1}$, then the degree of $p_{k}$ cannot be greater than $z+\lfloor(1+\varepsilon)z\rfloor$.
+By this logic, going from $A_{k-1}$ to $A_{k}$ consists of deleting the element of $A_{k-1}$ corresponding to the edges between a vertex in $S_{k-1}$ and $p_{k}$, which is $z$.
+Then, we will select $\lfloor(1+\varepsilon)z\rfloor$ elements of $A_{k-1}$ and increment them by $1$ (where we temporarily pad $A_{k-1}$ with zeroes to allow for the creation of new elements equal to 1).
+Example 9.
+Suppose that $A_{0}=\{3,2,2\}$, $\varepsilon=1/4$, and we wish to play a move on $z=4$. Then, $\lfloor(1+\varepsilon)z\rfloor=5$, and the possible multisets for $A_{1}$ are:
+$\{3,3,1,1,1\},\quad\{3,2,1,1,1,1\},\;\;\text{and}\;\;\{2,2,1,1,1,1,1\}.$
+The new multiset $A_{k}$ represents the edges between $p_{k}$ and a vertex in $V\setminus S_{k}$.
+Since any graph and sequence of best responses can be used to construct the multisets $(A_{k})_{0\leq k\leq m}$ in such a manner, it suffices to study the behavior of these moves.
 Definition 10.
-A move on a sequence $E_{k}$ consists of the following steps:
-  • Select some integer $z$ from $E_{k}$.
-  • Delete one instance of $z$ from $E_{k}$.
-  • Create $E_{k+1}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $E$ and adding 1 to each of them.
-For the rest of this proof, let $\sum E_{k}$ denote the sum of the elements of $E_{k}$.
+A regular move on a multiset $A_{k}$ with a positive integer $z\in A_{k}$ consists of the following steps:
+  • Delete one instance of $z$ from $A_{k}$.
+  • Pad $A_{k}$ with an sufficiently large (at least $z(1+\varepsilon)$) number of zeroes.
+  • Create $A_{k+1}$ from $A_{k}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $A_{k}$ and adding 1 to each of them.
+  • Remove all zeroes from $A_{k+1}$.
+As a shorthand notation, let $\sum A_{k}$ denote the sum of the elements (with their multiplicities) of $A_{k}$, and let $\sum A_{k}^{2}$ denote the sum of squares of the elements (with their multiplicities) of $A_{k}$.
+The former will be a useful measure to bound because of Lemma 12, and the latter will assist us in obtaining such a bound.
+Proposition 11.
+$\sum A_{0}=O\left(n|S_{0}|\right)$.
+Proof.
+Since there are only $n$ vertices in the graph, every vertex in $S_{0}$ can have at most $n$ edges.
+Recall that $\sum A_{0}$ counts the number of edges between a vertex in $S_{0}$ and a vertex in $V\setminus S_{0}$.
+Proposition 11 follows.
+∎
 Lemma 12.
-If $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$, then the final number of bad edges in the graph is $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ for all $\varepsilon=\omega\left(n^{-1/2}\right)$.
+If $\sum A_{m}=O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$, then the final number of bad edges in the graph is $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ if $\varepsilon=\omega(n^{-1/3})$.
 Proof.
-Let $T_{k}$ denote the number of edges that are incident to two vertices in $S_{k}$.
-Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum E_{k}$ denote $(\sum E_{k})-(\sum E_{k-1})$.
-For every move that we play on an element $z$ in $E_{k}$, the sum $\sum E_{k}$ grows by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges that are contained inside of $S_{k}$ grows by $z$.
+Let $T_{k}$ denote the number of edges between two vertices in $S_{k}$.
+Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum A_{k}$ denote $(\sum A_{k})-(\sum A_{k-1})$.
+For every regular move that we play on $A_{k-1}$ with an integer $z\in A_{k-1}$, the sum $\sum A_{k}$ grows from $\sum A_{k-1}$ by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges between two vertices in $S_{k}$ grows from $S_{k-1}$ by $z$.
 This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$.
 We have
-$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum E_{k}$
-$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum E_{k}$
-$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum E_{k}$
+$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum A_{k}$
+$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum A_{k}$
+$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum A_{k}$
 By summing both sides from $k=1$ to $m$, we get
-$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum E_{m}-\sum E_{0}.$
-This lemma requires that $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$.
-$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}m^{2}|S_{0}|\right)-\sum E_{0}$
+$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum A_{m}-\sum A_{0}.$
+Since $\sum A_{m}=O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$,
+$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}.$
 By Lemma 7, this becomes
-$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}$
-$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}+\Theta(n)+\varepsilon T_{0}$
-$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)+\Theta(n)+\varepsilon T_{0}$
-$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+\Theta\left(\varepsilon^{-1}n\right)+T_{0}.$
-Since $\Omega\left(\varepsilon^{-1}\right)$ bad edges are required to play the first uncertain best response, we must have $S_{0}=\Omega\left(\varepsilon^{-1}\right)$.
-This means that $O\left(\varepsilon n^{2}\right)$ must also be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$.
-If $\varepsilon=\omega\left(n^{-1/2}\right)$, then $O\left(\varepsilon n^{2}\right)$ will dominate $\Theta\left(\varepsilon^{-1}n\right)$.
+$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}$
+$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}+\Theta(n)+\varepsilon T_{0}$
+$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)-\varepsilon^{-1}\sum A_{0}+\Theta\left(\varepsilon^{-1}n\right)+T_{0}.$
+By Proposition 11, $\varepsilon^{-1}\sum A_{0}$ must be $O(\varepsilon^{-1}n|S_{0}|)$.
+Since $\varepsilon=\omega(n^{-1/3})$, this is dominated by the $O(\varepsilon^{2}n^{2}|S_{0}|)$ term.
+Also, since we must have $\Omega(\varepsilon^{-1})$ starting bad edges, $S_{0}$ is $\Omega(\varepsilon^{-1})$ as well.
+This means that $O(\varepsilon n^{2})$ must also be $O(\varepsilon^{2}n^{2}|S_{0}|)$.
+Finally, $\varepsilon=\omega(n^{-1/3})$, so $O(\varepsilon n^{2})$ will dominate $\Theta(\varepsilon^{-1}n)$.
+We are left with the following terms on the RHS:
 $T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+T_{0}.$
-Any vertex in $G$ is either in $S_{m}$, or $L$.
-This means that the only bad edges that are not counted by $T_{m}$ must be incident to a vertex in $S_{m}$ and a vertex in $L$.
-The edges that are incident to a vertex in $S_{m}$ and $G\setminus S_{m}=L$ are entirely described by $E_{m}$.
-Thus, the number of bad edges that exist between a vertex in $S_{m}$ and a vertex in $L$ is given by $\sum E_{m}$, which is $O\left(\varepsilon^{2}m^{2}\right)$.
-This is dominated by $T_{m}=O\left(\varepsilon n^{2}\right)+T_{0}$, so the asymptotic growth rate of the number of bad edges in the graph is given solely by that of $T_{m}$.
-It remains to show that $T_{0}$ cannot grow faster than $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)\geq\Theta(\varepsilon n^{2})$.
+We will now show that $T_{0}$ cannot grow faster than $O(\varepsilon^{2}n^{2}|S_{0}|)$.
 By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$.
-By Lemma 6, we start with $O\left(\varepsilon^{-2}\right)$ bad edges, so $T_{0}=O\left(\varepsilon^{-2}\right)$.
-This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega\left(n^{-1/2}\right)\geq\Theta\left(n^{-2/3}\right)$.
+By Proposition 5, we start with $O(\varepsilon^{-2})$ bad edges, so $T_{0}=O(\varepsilon^{-2})$.
+This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega(n^{-1/3})$.
+Then, $O(\varepsilon n^{2})$ is $O(\varepsilon^{2}n^{2}|S_{0}|)$ because $|S_{0}|=\Omega(\varepsilon^{-1})$.
+Our equation finally becomes
+$T_{m}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right).$
+Any vertex in $V$ is either in $S_{m}$, or $Q$.
+This means that the only bad edges that are not counted by $T_{m}$ must be between $S_{m}$ and a vertex in $Q$, which are entirely described by the elements in $A_{m}$.
+Thus, the number of bad edges between a vertex in $S_{m}$ and a vertex in $Q$ is bounded by $\sum A_{m}$, which is $O(\varepsilon^{3}n^{2}|S_{0}|)$.
+The total number of bad edges is then bounded by
+$T_{m}+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)=O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)=O\left(\varepsilon^{2}n^{2}|S_{0}|\right).$
+Intuitively, playing a regular move on $\sum A_{k-1}$ with a large $z$ would result in a big increase in from $\sum A_{k-1}$ to $\sum A_{k}$.
+However, playing such a move might also create many small elements in $A_{k}$.
+Having the sum $\sum A_{k}$ spread out across many small elements will make increasing $\sum A_{k}$ in future moves more difficult.
+We consider $\sum A_{k}^{2}$ as a way to measure how ”spread out” the multiset $A_{k}$ is, and this particular quantity is preferable over other measures because the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ can be related to $\sum A_{k-1}$.
+Our plan is to bound the behavior of $\sum A_{k}$ and $\sum A_{k}^{2}$ as $k$ increases. The multisets $A_{0}$ and $A_{m}$ will be subject to the following inequalities:
+$\sum A_{m}^{2}$ $\geq 0$ (1)
+$\sum A_{0}^{2}$ $\leq\left(\sum A_{0}\right)^{2}.$ (2)
+Eq.
+1 is true because the elements of $E_{m}$ must be nonnegative integers, and Eq.
+2 follows from the Cauchy-Schwarz inequality.
+We will only consider these inequalities (and not those with $m$ and 0 swapped) because we only need to prevent too many regular moves from being played with large $z$, which will naturally cause $\sum A_{k}^{2}$ to become small as $k$ increases.
+We will begin by investigating the behavior of the sum of squares.
+Let the current multiset be $A_{k-1}$.
+Then, pick some element $z\in A_{k-1}$, and define $\zeta=\lfloor z(1+\varepsilon)\rfloor$.
+We can play a regular move on $A_{k-1}$ with $z$, which will involve incrementing $\zeta$ the elements $a_{1},a_{2},\dots,a_{\zeta}\in A_{k-1}$ to $a_{1}+1,a_{2}+1,\dots,a_{\zeta}+1\in A_{k}$.
+It is easy to see that the effect of such a move on the increase from $\sum A_{k-1}$ to $\sum A_{k}$ is
+$\sum A_{k}-\sum A_{k-1}$ $=-z+\sum_{i=1}^{\zeta}\big((a_{i}+1)-a_{i}\big)$
+$\sum A_{k}-\sum A_{k-1}$ $=-z+\zeta.$ (3)
+Next, we must have
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+\sum_{i=1}^{\zeta}\left((a_{i}+1)^{2}+a_{i}^{2}\right)$
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta.$ (4)
+Eq.
+3 and Eq.
+4 together describe a single regular move.
+However, we can establish bounds upon these equations that are easier to analyze.
+Since $a_{1},a_{2},\dots,a_{\alpha}$ are all elements of $A_{k-1}$, their sum must be bounded by $\sum A_{k-1}$.
+By plugging this into our expression for the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$, we get
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta$
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum_{i=1}^{\zeta}a_{j}-z^{2}$
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum A_{k-1}-z^{2}.$ (5)
+Proposition 13.
+Taking the upper bound in Eq. 5 for the increase from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ never decreases $\sum A_{m}$.
+Proof.
+Notice that the changes described in the right hand sides of Eq.
+3, Eq.
+4, and Eq.
+5 are all independent of the current value of $\sum A_{k-1}^{2}$.
+Furthermore, the only constraint on the final sum of squares $\sum A_{m}^{2}$ is a lower bound provided by Eq.
+1.
+This means that always taking the upper bound for $\sum A_{k}$ from Eq.
+5 will never decrease $\sum A_{m}$.
 ∎
+Now equipped with Proposition 13, we actually no longer need that $z\in A_{k-1}$.
+In fact, we no longer even need to consider $A_{k-1}$ as a multiset anymore.
+For the rest of the proof, we will simply assume that $z$ is a positive real number.
+The next step is to find an upper bound for the change from $\sum A_{k-1}$ to $\sum A_{k}$.
+$\sum A_{k}-\sum A_{k-1}$ $=\zeta-z$
+$\sum A_{k}-\sum A_{k-1}$ $=\lfloor z(1+\varepsilon)\rfloor-z$
+$\sum A_{k}-\sum A_{k-1}$ $\leq z(1+\varepsilon)-z+1$
+$\sum A_{k}-\sum A_{k-1}$ $\leq\varepsilon z+1.$ (6)
+Proposition 14.
+Taking the upper bound in Eq. 6 for the increase from $\sum A_{k-1}$ to $\sum A_{k}$ never decreases $\sum A_{m}$.
+Proof.
+Suppose that we make a move using updates for the sum and sum of squares described by Eq.
+3 and Eq.
+5 respectively with some value for $z$.
+Then, we will show that there always exists another value $z^{\prime}$ with which we can make a move using Eq.
+6 and Eq.
+5 that results in the same value of $\sum A_{k}$ but a larger value of $\sum A_{k}^{2}$.
+By the same logic that was used in the proof of Proposition 13, this will never decrease $\sum A_{m}$.
+Since the values of $\sum A_{k}$ must match, our selection of $z^{\prime}$ must satisfy
+$\varepsilon z^{\prime}=\zeta-z=\lfloor z(1+\varepsilon)\rfloor-z.$
+The RHS of this equation is clearly positive, so $z^{\prime}$ must be nonpositive as well. We can also bound the size of $z^{\prime}$ with
+$\varepsilon z^{\prime}$ $=\lfloor z(1+\varepsilon)\rfloor-z$
+$\varepsilon z^{\prime}$ $\leq z(1+\varepsilon)-z$
+$z^{\prime}$ $\leq z.$
+By (5), smaller values of $z$ yield larger values of $\sum A_{k}^{2}$. This is exactly what we wanted. ∎
+With Proposition 13 and Proposition 14, we now have a strictly stronger set of actions on $\sum A_{k-1}$ and $\sum A_{k-1}^{2}$ that we can instead use with each of our moves.
+Definition 15.
+A strong move with $z$ on $(\sum A_{k-1},\sum A_{k-1}^{2})$ makes the following updates:
+$\sum A_{k}$ $=\sum A_{k-1}+\varepsilon z+1$
+$\sum A_{k}^{2}$ $=\sum A_{k-1}^{2}+2\sum A_{k-1}-z^{2}.$
+For all integers $1\leq k\leq m$, let $z_{k}$ denote the value of $z$ used in the strong reversed move from $(\sum A_{k-1},\sum A_{k-1}^{2})$ to $(\sum A_{k},\sum A_{k}^{2})$. Then,
+$\sum A_{k}$ $=\sum A_{0}+k+\varepsilon\sum_{i=1}^{k}z_{i}.$ (7)
+The same process of expanding is more complicated for $\sum A_{m}^{2}$:
+$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{i}\right)-\sum_{i=1}^{m}z_{i}^{2}.$
+We can rewrite the $\sum A_{j}$ by plugging in Eq. 1 and Eq. 7.
+$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{0}+i+\varepsilon\sum_{j=1}^{i}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
+$0$ $\leq\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\varepsilon\sum_{j=1}^{m}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
+$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum_{i=1}^{m}z_{i}^{2}$ (8)
+By the Cauchy-Schwarz Inequality,
+$\left(\sum_{j=1}^{m}1\right)\left(\sum_{j=1}^{m}z_{j}^{2}\right)$ $\geq\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
+$\sum_{j=1}^{m}z_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}.$
+Plugging this result into Eq. 8 yields a quadratic polynomial in terms of $\sum_{i=1}^{m}z_{i}$ in the RHS.
+$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
+$0$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}-2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum A_{0}^{2}$
+Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}z_{j}$ must be less than the larger root of this quadratic.
+This larger root is obtained through a tedious application of the quadratic formula:
+$\sum_{j=1}^{m}z_{j}\leq\frac{2\varepsilon m+\sqrt{4\varepsilon^{2}m^{2}+4m^{-1}\sum A_{0}^{2}}}{2m^{-1}}$ (9)
 Lemma 16.
-The quantity $\sum E_{0}$ grows on the order of $O\left(m|S_{0}|\right)$.
-Proof.
-Since there are only $n$ vertices in the graph, every vertex in $S_{0}$ can have at most $n$ edges.
-Recall that $\sum E_{0}$ counts the number of edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$.
-By Lemma 8, we have
-$\sum E_{0}$ $=O\left(n|S_{0}|\right)$
-$\sum E_{0}$ $=O\left(m|S_{0}|\right).$
-We will now consider the time reversed process of playing moves on the $E_{k}$.
-Definition 12.
-A reversed move on a sequence $E_{k}$ consists of the following steps:
-  • Select a list of positive elements $\{e_{1},e_{2},\dots,e_{\alpha}\}$ from $E_{k}$
-  • Subtract one from each of the $e_{j}$ for $1\leq j\leq\alpha$
-  • Calculate $z=\lceil\alpha/(1+\epsilon)\rceil$
-  • Create $E_{k-1}$ by inserting $z$ into $E_{k}$ as a new element
-Let $\sum E_{k}^{2}$ denote the sum of the squares of the elements of $E_{k}$.
-In order to investigate the asymptotic growth rate of $\sum E_{m}$, we wish to find the largest starting value of $\sum E_{m}$ for which it is possible to play $m$ reversed moves to get to a desired value of $\sum E_{0}$.
-Our plan is to bound the behavior of $\sum E_{k}$ and $\sum E_{k}^{2}$ as reversed moves are played.
-Intuitively, playing a reversed move on many elements of $E_{k}$ would result in a large decrease in from $\sum E_{k}$ to $\sum E_{k-1}$.
-However, playing such a move would also result in the creation of a new, very large element in $E_{k-1}$.
-Since we can only ever subtract one from each element per move, having the sum $\sum E_{k-1}$ concentrated in a few large elements will make decreasing $\sum E$ more difficult.
-We consider $\sum E_{k}^{2}$ as a way to measure how ”spread out” the sequence $E_{k}$ is, and this particular quantity is preferable over other measures because of how the change from $\sum E_{k}^{2}$ to $\sum E_{k-1}^{2}$ can be cleanly bounded by an expression in terms of $\sum E_{k}$.
-Consider some arbitrary sequence $E_{k}$.
-Let $\alpha$ denote the number of elements of $E$ that we play the current reversed game move on, and let these elements be denoted $e_{1},e_{2},\dots,e_{\alpha}$.
-By the rules of the reversed game, playing such a move will create a new element of size $z=\lceil\alpha/(1+\varepsilon)\rceil$.
-It is easy to see that the effect of such a move on $\sum E_{k}$ is
-$\sum E_{k-1}=\sum E_{k}-\alpha+z.$
-Next, we must have
-$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-\sum_{j=1}^{\alpha}\left(e_{j}^{2}-(e_{j}-1)^{2}\right)+z^{2}$
-$=\sum E_{k}^{2}-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}.$
-Since $e_{1},e_{2},\dots,e_{\alpha}$ are all elements of $E_{k}$, their sum must be bounded by $\sum E_{k}$. By plugging this into our expression for the change in $\sum E^{2}$, we get
-$\sum E_{k-1}^{2}-\sum E_{k}^{2}$
-$={}$ $-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}$
-$\geq{}$ $-2\sum_{j=1}^{\alpha}e_{j}+z^{2}$
-$\geq{}$ $-2\sum E_{k}+\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil^{2}$
-$\geq{}$ $-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}$ (1)
-Lemma 13.
-It is always optimal to take the lower bound in (1) for the change in $\sum E^{2}$.
-Proof.
-Notice that neither our change in $\sum E$ nor our bound (1) on the change in $\sum E^{2}$ depend on $\sum E_{k}^{2}$.
-If we make $\sum E_{k}^{2}$ as small as possible for a fixed $\sum E_{k}$, the future options we will have for $\alpha$ will not change.
-Furthermore, we only care that $\sum E_{0}^{2}$ ends up being smaller than $\left(\sum E_{0}\right)^{2}$.
-This means that allowing $\sum E_{k-1}^{2}$ to attain its lower bound will never result in more moves being required.
-There are no further issues because there is no need that intermediate values can be represented by actual sequences; there is no required minimum value for $\sum E_{k}^{2}$ for any integer $0\leq k\leq m-1$.
-∎
-We can also obtain a simpler lower bound for the change in $\sum E$.
-$\sum E_{k-1}-\sum E_{k}$
-$={}$ $-\alpha+z$
-$={}$ $\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
-$\geq{}$ $(1-\varepsilon)\alpha-\alpha$
-$={}$ $-\varepsilon\alpha.$ (2)
-Lemma 14.
-It is always optimal to take the lower bound in (2) for the change in $\sum E$.
-Proof.
-If we allow $\alpha\in\mathbb{R}$, still with $\alpha\geq 1$, it is always possible to choose a new $\alpha^{\prime}$ such that
-$\varepsilon\alpha^{\prime}=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha.$
-The RHS of the previous equation is clearly nonpositive, so $\alpha^{\prime}$ must be nonpositive as well. We can also bound the size of $\alpha^{\prime}$ with
-$-\varepsilon\alpha^{\prime}$ $=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
-$1-\varepsilon\alpha^{\prime}$ $\geq 1+\frac{\alpha}{1+\varepsilon}-\alpha$
-$-\varepsilon\alpha^{\prime}$ $\geq\frac{-\varepsilon\alpha}{1+\varepsilon}$
-$(1+\varepsilon)\alpha^{\prime}$ $\leq\alpha$
-$\alpha^{\prime}$ $<\alpha.$
-By (1), if we take the lower bound for the change in $\sum E_{k}^{2}$ using $\alpha^{\prime}$, the value of $\sum E_{k-1}^{2}$ is equal to or smaller than if we had used $\alpha$ instead.
-By the same logic that was used to prove Lemma 13, it is always helpful to have a lower $\sum E^{2}$.
-∎
-With Lemma 13 and Lemma 14, we now have a strictly stronger set of actions on $\sum E$ and $\sum E^{2}$ that we can perform with each move.
-Definition 15.
-A strong reversed move that uses $\alpha$ on the pair of values $\sum E_{k}$ and $\sum E_{k}^{2}$ is given by
-$\sum E_{k-1}$ $=\sum E_{k}-\varepsilon\alpha$
-$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}.$
-For integers $1\leq k\leq m$, let $\alpha_{k}$ denote the value of $\alpha$ used in the strong reversed move to get from $\left(\sum E_{k},\sum E_{k}^{2}\right)$ to $\left(\sum E_{k-1},\sum E_{k-1}^{2}\right)$.
-Then,
-$\sum E_{m}$ $=\sum E_{0}+\varepsilon\alpha_{1}+\varepsilon\alpha_{2}+\cdots+\varepsilon\alpha_{m}$
-$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}.$ (3)
-The same process is a little more complicated for $\sum E_{m}^{2}$.
-$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-We can rewrite the $\sum E_{j}$ by plugging in (3).
-$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{0}+\varepsilon\sum_{i=1}^{j}\alpha_{i}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2m\sum E_{0}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-$0$ $=\sum E_{0}^{2}+2m\sum E_{0}-\sum E_{m}^{2}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-We have $\sum E_{m}^{2}\geq\sum E_{m}$. Furthermore, due to (3), $\sum E_{m}\geq\varepsilon\sum_{j=1}^{m}\alpha_{j}$. We can use these two inequalities to replace $\sum E_{m}^{2}$.
-$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
-$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$ (4)
-By the Cauchy-Schwarz Inequality, we have
-$\left(\sum_{j=1}^{m}1\right)\left(\sum_{j=1}^{m}\alpha_{j}^{2}\right)$ $\geq\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
-$\sum_{j=1}^{m}\alpha_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}.$
-We can then plug this into (4):
-$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
-$0$ $\geq(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}-\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-\left(\sum E_{0}^{2}+2m\sum E_{0}\right).$
-We are now left with a quadratic in the sum $\sum_{j=1}^{m}\alpha_{j}$.
-Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}\alpha_{j}$ must be less than the larger root of this quadratic.
-This larger root is obtained through a tedious application of the quadratic formula:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\varepsilon(2m-1)+\sqrt{\varepsilon^{2}(2m-1)^{2}+4(1-\varepsilon)^{2}m^{-1}\left(\sum E_{0}^{2}+2m\sum E_{0}\right)}}{(1-\varepsilon)^{2}m^{-1}}.$ (5)
-Lemma 16.
-If $\varepsilon=\Omega\left(n^{-1/4}\right)$, then $\sum_{j=1}^{m}\alpha_{j}$ is bounded by $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
+Suppose that $\varepsilon=\Omega(n^{-1/4})$. Then, $\sum_{i=1}^{m}z_{j}$ is $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
 Proof.
 If we substitute asymptotic behavior into the RHS of (9), we get
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum E_{0}^{2}\right)+\Theta\left(\sum E_{0}\right)}}{\Theta\left(m^{-1}\right)}.$
-Notice that we must have $\sum E_{0}^{2}\leq\left(\sum E_{0}\right)$. By Lemma 11, we can rewrite the previous inequality as follows:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)+\Theta\left(m|S_{0}|\right)}}{\Theta\left(m^{-1}\right)}$
-It is clear that the $\Theta\left(m|S_{0}|^{2}\right)$ term dominates the $\Theta\left(m|S_{0}|\right)$ term. We now have the following:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}.$ (6)
+$\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum A_{0}^{2}\right)}}{\Theta\left(m^{-1}\right)}.$
+By Lemma 7, we can replace every $m$ in the RHS with an $n$ to get
+$\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+\Theta\left(n^{-1}\sum A_{0}^{2}\right)}}{\Theta\left(n^{-1}\right)}.$
+By Eq. 2 and Proposition 11, $\sum A_{0}^{2}\leq(\sum A_{0})^{2}=O(n^{2}|S_{0}|^{2})$, so we can rewrite the previous inequality as follows:
+$\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}.$ (10)
 Depending on which one of the two terms inside of the square root dominate, there are two cases that we have to tackle.
 Case 16.1.
-$\Theta\left(\varepsilon^{2}m^{2}\right)$ dominates $\Theta\left(m|S_{0}|^{2}\right)$.
-In this case, we can rewrite (6) as follows:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)}}{\Theta\left(m^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)}{\Theta\left(m^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon m^{2}\right).$
-Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
+$\Theta\left(\varepsilon^{2}n^{2}\right)$ dominates $\Theta\left(n|S_{0}|^{2}\right)$.
+In this case, we can rewrite Eq. 10 as follows:
+$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon n^{2}\right).$
+Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$ because we must begin with that many bad edges.
+This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
 Case 16.2.
 $\Theta\left(m|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}m^{2}\right)$.
-In this case, we can rewrite (6) as follows:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(m^{3/2}|S_{0}|\right).$
-By Lemma 8, $\varepsilon=\Omega\left(n^{-1/4}\right)$ is equivalent to $\varepsilon=\Omega\left(m^{-1/4}\right)$.
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}m^{3/2}|S_{0}|\right)$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(m^{-1/4}\right)^{-2}m^{3/2}|S_{0}|\right)$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$
-The result of Lemma 16 can be plugged into (3) with Lemma 11 to obtain a bound on $\sum E_{m}$.
-Since we are only interested in $\varepsilon=\Omega\left(n^{-1/4}\right)\geq\Theta\left(n^{-1/3}\right)$, we have
-$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}$
-$\sum E_{m}$ $=O\left(m|S_{0}|\right)+O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$
-$\sum E_{m}$ $=O\left(\varepsilon^{3}m^{2}|S_{0}|\right).$ (7)
-By (7) and Lemma 10, the final number of bad edges in the graph must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$.
+In this case, we can rewrite Eq. 10 as follows:
+$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(n^{3/2}|S_{0}|\right).$
+Since $\varepsilon=\Omega(n^{-1/4})$,
+$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}n^{3/2}|S_{0}|\right)$
+$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(n^{-1/4}\right)^{-2}n^{3/2}|S_{0}|\right)$
+$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$
+The result of Lemma 16 can be plugged into Eq. 7 with Proposition 11 to obtain a bound on $\sum E_{m}$. Since $\varepsilon=\Omega(n^{-1/4})$, we have
+$\sum A_{m}$ $=\sum A_{0}+\varepsilon\sum_{i=1}^{m}z_{i}$
+$\sum A_{m}$ $=O\left(n|S_{0}|\right)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$
+$\sum A_{m}$ $=O\left(\varepsilon^{3}n^{2}|S_{0}|\right).$ (11)
+By Eq.
+11 and Lemma 12, the final number of bad edges in the graph must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$.
 By construction, every vertex in $S_{0}$ begins with at least one bad edge, so we must begin with $\Omega\left(|S_{0}|\right)$ bad edges.
-We obtain the following upper bound on the price of uncertainty for consensus games when $\varepsilon=\Omega\left(n^{-1/4}\right)$:
-$\frac{O\left(\varepsilon^{2}n^{2}|S_{0}|\right)}{\Omega\left(|S_{0}|\right)}=O\left(\varepsilon^{2}n^{2}\right).$
+We obtain the following upper bound on the price of uncertainty for consensus games when $\varepsilon=\Omega(n^{-1/4})$:
+$\operatorname{PoU}(\varepsilon,\text{consensus})=\frac{O\left(\varepsilon^{2}n^{2}|S_{0}|\right)}{\Omega\left(|S_{0}|\right)}=O\left(\varepsilon^{2}n^{2}\right).$
 This finishes the proof of Theorem 4.
+5 Conclusion
+In this paper, we established tight asymptotic bounds on the worst-case increase in social cost experienced by a network under uncertainty.
+Our work leaves open some interesting future research directions.
+Firstly, we believe that the lower bound also holds in the more general case of arbitrary agent update ordering.
+This may more closely reflect the dynamics of real social networks, and we have a modified version of our construction in this paper that could potentially work for this new problem.
+Additionally, the study of adversarial influences to normal operations due to uncertainty or individual biases could also be extended to other game-theoretic network settings.
+While agents in our consensus game model have fixed relationships defined by graph edges, our model does not totally capture the fluidity of social networks due to the formation and destruction of individual connections.
+Incorporating uncertainty into more general models for social networks can provide greater insight into their behavior in the real world.

### v3

@@
 Modeling uncertainty as perturbations of relative magnitude $1+\varepsilon$ to these color neighbor counts, we show that even small values of $\varepsilon$ greatly hinder the ability to achieve consensus in a social network.
-We prove theoretically tight upper and lower bounds on the price of uncertainty, a metric defined by Balcan et al.
-(2009) to quantify the effect of uncertainty in network games.
+We prove theoretically tight upper and lower bounds on the price of uncertainty, a metric defined in previous work by Balcan et al.
+to quantify the effect of uncertainty in network games.
 1 Introduction
-In the past decade the rapid expansion of social media platforms has resulted in the formation of large and highly connected social networks.
+In the past decade, the rapid expansion of social media platforms has resulted in the formation of large and highly connected social networks.
 The evolution of these networks has been extensively studied in order to better understand how opinions held by agents are affected by their interactions with other people.
-In a social setting, individuals generally tend to modify their beliefs to more closely align with others that they frequently interact with (Moussaïd et al., 2013; Garimella et al., 2018).
-This generally causes opinions held by agents in a network to generally converge over time as interacting agents attempt to reach a consensus.
-However, previous research has shown that additional factors, such as a sparse network structure or agents who refuse to change their opinion, can hinder consensus formation.
-(Acemoğlu et al., 2013; Hegselmann and Krause, 2002; Moussaïd et al., 2013).
+In a social setting, individuals often modify their beliefs to more closely align with others that they frequently interact with (Moussaïd et al., 2013; Garimella et al., 2018).
+This generally causes opinions held by agents in a network to converge over time as interacting agents attempt to reach a consensus.
+However, previous research has shown that additional factors, such as a sparse network structure or agents who refuse to change their opinions, can hinder consensus formation (Acemoğlu et al., 2013; Hegselmann and Krause, 2002; Moussaïd et al., 2013).
 The vast majority of social network models in the existing literature assume that agents have accurate information about the opinions held by others.
-However, in large networks where vertices generally have many neighbors and where interactions between agents are often brief, agents are often unable to make accurate judgments about the opinions of other agents such as the number of their consenting or dissenting peers (Xiong and Liu, 2014).
+However, in large networks where vertices generally have many neighbors and where interactions between agents are often brief, agents are often unable to make accurate judgments about the opinions of other agents, such as the number of their consenting or dissenting peers (Xiong and Liu, 2014).
 Previous work modeling uncertainty in games has focused on the failure of agents to always pick the action with the optimal payoff.
-For example, in the quantal response model, agents sometimes deviate from optimal behavior with strategies providing higher payoffs being chosen more often than those providing lower payoffs (Penna, 2018; Blume, 1993; Alós-Ferrer and Netzer, 2010).
+For example, in the quantal response model, agents sometimes deviate from optimal behavior, with strategies providing higher payoffs being chosen more often than those providing lower payoffs (Penna, 2018; Blume, 1993; Alós-Ferrer and Netzer, 2010).
 Alternatively, agents may occasionally fail to participate in the game, which can sometimes increase the social cost (Meir et al., 2012; Li et al., 2017).
@@
 We consider a model of uncertainty first proposed by Balcan et al.
-(2009) where agents experience small, adversarial perturbations of magnitude $1+\varepsilon$ to their observations of their neighbors.
+(2009), where agents experience small, adversarial perturbations of relative magnitude $1+\varepsilon$ to their observations of their neighbors.
 These perturbations capture the difficulty of ascertaining accurate information in large social networks, as the absolute margin of error increases proportionally to the number of neighbors an agent has.
@@
 They quantify the effect of uncertainty using a measure called the price of uncertainty (PoU) of consensus games, defined as the maximum multiplicative factor by which the social cost can increase from any starting state.
-Even extremely small perturbations can allow adversarial perturbations to snowball out of control and cause the game to move to states of much higher social cost (Balcan et al., 2011).
+Even extremely small adversarial perturbations can snowball out of control and cause the game to move to states of much higher social cost (Balcan et al., 2011).
 Our main results are tight asymptotic upper and lower bounds on the PoU for consensus games.
@@
 In this paper, we extend their work by proving that the PoU for consensus games is exactly $\Theta(n^{2}\varepsilon^{2})$.
-We believe that our result also holds irrespective of agent update ordering (i.e., the adversary can only control the magnitude of the perturbations), and this a potential future direction for research.
+We believe that our result also holds irrespective of agent update ordering (i.e., the adversary can only control the magnitude of the perturbations), and this is a potential future direction for research.
 2 The Model
-We consider a consensus game, where each of $n$ agents are located at distinct vertices of a simple undirected graph with $n$ vertices (Christodoulou et al., 2012).
+We consider a consensus game in which each of $n$ agents is located at a distinct vertex of a simple undirected graph with $n$ vertices (Christodoulou et al., 2012).
 In this paper, we will use the words agent and vertex interchangeably.
 Each agent can choose a color from $\{0,1\}$.
-Let us refer to two agents as neighbors iff there exists an edge between the two agents.
-An agent’s cost is defined as the number of their neighbors that pick a different color, which we call bad neighbors.
-We will also call edges whose endpoints have different colors bad edges.
-Finally, for any state $S$ of the game, we define a social cost function $\operatorname{cost}(S)$ as the sum of the costs of every agent in the game.
+Let us refer to two agents as neighbors if and only if there exists an edge between them.
+An agent’s cost is defined as the number of its neighbors that pick a different color; we call these bad neighbors.
+We will also call edges whose endpoints have different colors bad edges, and we will call neighbors or edges that are not bad good.
+Finally, for any state $S$ of the game, we define a social cost function $\operatorname{cost}(S)$ as the sum of the costs of all agents in the game.
 Starting from a ground state $S_{0}$ with social cost at least 1, let $S_{t}$ be a state achieved after $t$ steps under uncertain best-response dynamics.
-Here, the term uncertain here refers to the fact that an agent might miscount their cost by a multiplicative factor of at most $1+\varepsilon$ for an uncertainty factor $\varepsilon>0$ that can depend on $n$.
+Here, the term uncertain refers to the fact that an agent might miscount its cost by a multiplicative factor of at most $1+\varepsilon$ for an uncertainty factor $0<\varepsilon<1$ that can depend on $n$.
 Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty.
-By definition, an uncertain best response must involve a color switch.
+In particular, the number of bad neighbors of the agent playing an uncertain best response must not exceed the number of good neighbors by more than a multiplicative factor of $1+\varepsilon$.
+Also, by definition, an uncertain best response must involve a color switch.
 Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise an agent cannot miscount its number of bad neighbors and hence there is no uncertainty.
@@
 where the maximum is taken over all initial states $S_{0}$ and all states $S_{t}$ reachable from $S_{0}$ in some number $t$ of steps (Balcan et al., 2009).
-This paper obtains a lower bound for the PoU of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
+This paper obtains a lower bound of $\Omega\left(\varepsilon^{2}n^{2}\right)$ on the PoU of consensus games for all $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
 Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
-This is achieved using an intricate construction that amplifies the “snowball” effect of uncertainty in Balcan et al.
+This is achieved using an intricate construction that amplifies the “snowball” effect of uncertainty described in Balcan et al.
 (2011).
-Next, we prove that the PoU for consensus games is also bounded from above by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega(\varepsilon^{-1/4})$ by analyzing a clever monovariant.
+Next, by analyzing a clever construction, we prove that the PoU for consensus games is also bounded from above by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega(n^{-1/4})$.
 These bounds coincide, so we conclude that the asymptotically tight PoU for consensus games is $\Theta(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$, resolving an open problem of Balcan et al.
@@
 (2011) utilizes an initializer gadget to set up initial colorings for an output component.
-This component consists of successive layers of vertices, where each layer has more vertices than the previous layer by a factor of $(1+\varepsilon)$, and every vertex is connected to all vertices on adjacent layers.
+This component consists of successive layers of vertices, each of which has more vertices than the previous layer by a factor of $(1+\varepsilon)$, and every vertex is connected to all vertices on adjacent layers.
 This output component is responsible for generating $\Omega(\varepsilon n^{2})$ bad edges, and the initializer gadget requires $\Theta(1/\varepsilon^{2})$ starting bad edges, so this achieves a lower bound of $\Omega(\varepsilon^{3}n^{2})$.
@@
 Upper Bound.
-Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process.
+Generally, the goal of an adversary that wants to blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best response process.
 Consider a time in which there are $m^{2}$ bad edges.
-The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$.
+The most “compact” way to store these bad edges using the least number of vertices is a bipartite graph $K_{m,m}$.
 If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$.
-However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response.
-Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
-For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$.
-Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again.
-Assuming most edges can eventually be turned bad again as is the case in our lower bound construction, the total number of bad edges at the end is
+However, this requires $(1+\varepsilon)m$ new vertices to “absorb” the new bad edges produced by the uncertain best-response.
+Hence, increasing the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
+Following this heuristic for a constant $\varepsilon$ (i.e., not depending on $n$), the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$.
+Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain best response, can be turned bad again.
+Assuming most edges can eventually be turned bad again, as is the case in our lower bound construction, the total number of bad edges at the end is
 $O\left(\varepsilon^{2}n^{2}+\frac{\varepsilon^{2}n^{2}}{1+\varepsilon}+\cdots\right)=O(\varepsilon n^{2}).$
-Since there must be $\Omega(1/\varepsilon)$ bad edges at the start, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
-It appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses.
+Since there must be $\Omega(1/\varepsilon)$ bad edges at the start, we estimate the upper bound on the price of uncertainty for consensus games to be $O(\varepsilon^{2}n^{2})$.
+It appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain best responses.
 The hope is that we can prove an upper bound on the growth rate in line with our estimate of $O(\varepsilon^{2}n^{2})$.
@@
 Example 1.
-Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega(1/\varepsilon)$.
-The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$.
-We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
-  • Let $V$ play an uncertain best response that switches its color to red, resulting in $m+1$ bad edges.
-  • Take the endpoint of one such bad edge adjacent to $W$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $W$.
+Consider a red vertex $V$ whose only neighbors are $m$ blue vertices and $m+1$ red vertices for $m=\Omega(1/\varepsilon)$.
+The $m$ blue vertices are all connected to another blue vertex $B$, and the $m+1$ red vertices are all connected to another red vertex $R$.
+Then, we can have $V$ play $2m+1$ uncertain best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices.
+  • Let $V$ play an uncertain best response that switches its color to blue, resulting in $m+1$ bad edges.
+  • Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $R$.
   • Switch the color of $V$ again, resulting in $m+1$ bad edges.
-Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $W$, and 1 bad edge incident to $R$.
-  • Repeat the above, at each step switching the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once.
-There are now $m$ bad edges incident to $R$ and $m+1$ bad edges incident to $W$.
-Note that Example 1 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges.
+Take the endpoint of one such bad edge adjacent to $B$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $B$, and 1 bad edge incident to $R$.
+  • Repeat the above, switching at each step the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once.
+There are now $m$ bad edges incident to $B$ and $m+1$ bad edges incident to $R$.
+Note that Example 1 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is blue, $m$ of the vertices in the $(2m+1)$-size part are red, and $m+1$ of the vertices in the $(2m+1)$-size part are blue to also roughly double the number of bad edges.
 The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in our initial estimate it was $1+\varepsilon$.
-Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step.
-We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
-Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the step at which each vertex first plays an uncertain best response.
+Hence, using this analysis, it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step.
+We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ on the number of bad edges at the last stage using this strategy.
+Instead, one idea is that we should focus not on the total number of bad edges at every step but rather consider the step at which each vertex first plays an uncertain best response.
 This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last.
@@
 This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
-Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” as a clever monovariant.
+Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” via a clever construction.
 3 Lower Bound
-We first improve the lower bound for the PoU for consensus games.
+We first improve the lower bound on the PoU for consensus games.
 Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
-Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
+Here, $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
 Theorem 2.
@@
 3.1 Proof of Improved Lower Bound
-One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level.
-Starting from the first level, there are enough bad edges for all vertices in the first level to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$.
+One general idea is to construct a graph as a sequence of levels, where each level contains slightly more vertices than the previous level.
+At the beginning, all of the vertices in the first level should have enough bad edges to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$.
 The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
-We now make the above idea rigorous.
+We will make the above idea rigorous.
 We define a boosting operation between layers of the following form.
@@
 Suppose that every vertex in level $\ell$ is connected to every vertex in level $\ell-1$ and level $\ell+1$, so the degree of every vertex in level $\ell$ is $d_{\ell}=\left(\frac{1}{1+\frac{\varepsilon}{2}}+(1+\frac{\varepsilon}{2})\right)\cdot m$.
-If the vertices in levels $\ell$ and $\ell+1$ are all white and the vertices in level $\ell-1$ are all red, then the edges between levels $\ell-1$ and $\ell$ are all bad.
-Since $(1+\varepsilon)\cdot\frac{1}{1+\frac{\varepsilon}{2}}\geq\frac{d_{\ell}}{2}$ for all $\varepsilon\leq 1$ then all vertices in level $\ell$ are allowed to play uncertain best responses in sequence, turning level $\ell$ all red.
+If the vertices in levels $\ell$ and $\ell+1$ are all blue and the vertices in level $\ell-1$ are all red, then the edges between levels $\ell-1$ and $\ell$ are all bad.
+Since $(1+\varepsilon)\cdot\frac{1}{1+\frac{\varepsilon}{2}}\geq\frac{d_{\ell}}{2}$ for all $\varepsilon\leq 1$, then all vertices in level $\ell$ are allowed to play uncertain best responses in sequence.
+This turns level $\ell$ all red.
 Note now that the number of bad edges between levels $\ell$ and $\ell+1$ is $\left(1+\frac{\varepsilon}{2}\right)\cdot m^{2}$, which is a $1+O(\varepsilon)$ multiplicative increase from the number of original bad edges $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m^{2}$ from before this boosting operation.
 We now present the construction.
-In order to implement the idea above, we must first find some way to create the first two layers of size $\frac{1}{\varepsilon}$ since a vertex must have degree at least $\frac{1}{\varepsilon}$ in order to play an uncertain best response.
+In order to implement the idea above, we must first find some way to create the first two layers of size $\frac{1}{\varepsilon}$ because a vertex must have degree at least $\frac{1}{\varepsilon}$ in order to play an uncertain best response.
 The initializer layers will have numbers of vertices
-$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{1}{\varepsilon},\frac{1}{\varepsilon}$
-where the odd indexed layers cycle the consecutive integers from $\frac{1}{\varepsilon}$ to $\frac{2}{\varepsilon}$ and the numbers of vertices in even indexed layers increase by a factor of 2 whenever the odd indexed layer’s cycle repeats until the number reaches $\frac{1}{\varepsilon}$.
-Each vertex is connected to all vertices in the layer immediately after it except for vertices in odd indexed layers with $\frac{2}{\varepsilon}$ vertices, which are connected to the next layer to form a regular bipartite graph with the next layer where the vertices in the next layer are each connected to $\frac{1}{\varepsilon}$ vertices in the current layer.
-Note that the total number of initializer layers is $\frac{1}{\varepsilon}\cdot\log_{2}\frac{1}{\varepsilon}$ and each layer has at most $\frac{2}{\varepsilon}$ vertices, so the total number of vertices in the initializer layers is $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$.
-The initializer sequence has been constructed in such a way that that starting from a state in which the first layer is red and all other layers are white, the vertices in each layer can play uncertain best responses to change their color to red in sequence so that the last two layers form a $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ where the left side is red and the right half is white.
-Finally, we construct $\frac{1}{\varepsilon}$ white vertices in a parallel layer to the red left side of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ and connect each of them to every vertex in the right half of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$, so that each vertex in the right half now has $\frac{1}{\varepsilon}$ red neighbors and $\frac{1}{\varepsilon}$ white neighbors on the left.
+$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{1}{\varepsilon},\frac{1}{\varepsilon},$
+where the numbers of vertices in the odd-indexed layers cycle through the consecutive integers from $\frac{1}{\varepsilon}$ to $\frac{2}{\varepsilon}$, and the numbers of vertices in even-indexed layers increase by a factor of 2 whenever the odd-indexed layers’ cycle repeats up until the number reaches $\frac{1}{\varepsilon}$.
+Each vertex is connected to all vertices in the layer immediately after it except for vertices in odd-indexed layers with $\frac{2}{\varepsilon}$ vertices, which are connected to the next layer to form a regular bipartite graph with the next layer where the vertices in the next layer are each connected to $\frac{1}{\varepsilon}$ vertices in the current layer.
+Note that the total number of initializer layers is $\frac{1}{\varepsilon}\cdot\log_{2}\frac{1}{\varepsilon}$, and each layer has at most $\frac{2}{\varepsilon}$ vertices, so the total number of vertices in the initializer layers is $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$.
+The initializer sequence has been constructed in such a way such that by starting from a state in which the first layer is red and all other layers are blue, the vertices in each layer can play uncertain best responses to change their color to red in sequence so that the last two layers form a $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ where the left side is red and the right half is blue.
+Finally, we construct $\frac{1}{\varepsilon}$ blue vertices in a parallel layer to the red left side of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ and connect each of them to every vertex in the right half of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$, so that each vertex in the right half now has $\frac{1}{\varepsilon}$ red neighbors and $\frac{1}{\varepsilon}$ blue neighbors on the left.
 Note that no new bad edges are added.
-We will now leave all layers before the current layer frozen, that is, no layer in a layer before the current layer will ever change color again.
-After the current layer, construct a secondary set of layers with numbers of vertices
-$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{2}{\varepsilon},\frac{2}{\varepsilon}.$
-Note that this sequence is almost identical to that of the initializer except we keep going until a $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ is formed by the last two layers.
-The first layer in the secondary set is the same as the last layer of the initializer, each vertex of which is connected with $\frac{1}{\varepsilon}$ frozen white vertices and $\frac{1}{\varepsilon}$ frozen red vertices.
-Note the following observation:
+We will now leave all layers before the last layer frozen: no vertex in a layer before the last layer will ever change color again.
+After the last layer in the initializer, construct a secondary set of layers with numbers of vertices
+$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{2}{\varepsilon},\frac{2}{\varepsilon}.$
+Note that this sequence is almost identical to that of the initializer, except we keep going until a $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ is formed by the last two layers.
+The first layer in the secondary set is the same as the last layer of the initializer, each vertex of which is connected with $\frac{1}{\varepsilon}$ frozen blue vertices and $\frac{1}{\varepsilon}$ frozen red vertices.
+We now have the following observation:
 Proposition 3.
@@
 ∎
-It is our careful construction of the initializer to get to this point combined with Proposition 3 that is the key observation needed to improve the bound in Balcan et al.
+Our careful construction of the initializer and Proposition 3 are the key observations needed to improve the bound in Balcan et al.
 (2011).
-We now finish the construction.
+We will now finish the construction.
 Note that the $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ produced at the end of the secondary set is a starting point from which we can perform the boosting operations as defined above.
-Here we add a layer of $\frac{2}{\varepsilon}+1$ vertices to form an initial $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}+1}$ and then initialize $m$ to be $\frac{2}{\varepsilon}+1$.
-Starting from this layer, layers will increase in size by a constant multiplicative factor of $1+\frac{\varepsilon}{2}$ until we reach $c\cdot\varepsilon n$ vertices in the last layer for some small constant $c$ to be set later.
-The total number of vertices used in the boosting operations is $\Theta\left(\frac{c\cdot\varepsilon n}{1-\frac{1}{1+\frac{\varepsilon}{2}}}\right)=\Theta(cn)$.
-Combined with the number of vertices $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$ used in the initializer as well as the secondary set, we see that for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$ and a constant $c$ independent of $\varepsilon$ and $n$ that the total number of vertices used is at most $n$, which means our construction has the right number of vertices.
+Here we add a layer of $\frac{2}{\varepsilon}+1$ vertices to form an initial $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}+1}$, and then initialize $m$ to be $\frac{2}{\varepsilon}+1$.
+Starting from this layer, additional layers will increase in size by a constant multiplicative factor of $1+\frac{\varepsilon}{2}$ until we reach $c\cdot\varepsilon n$ vertices in the last layer for some small constant $c$ to be set later.
+The total number of vertices used in the boosting operations is
+$\Theta\left(\frac{c\cdot\varepsilon n}{1-\frac{1}{1+\frac{\varepsilon}{2}}}\right)=\Theta\left(\frac{c\cdot\varepsilon n}{\frac{\varepsilon}{2+\varepsilon}}\right)=\Theta(cn).$
+Combined with the number of vertices $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$ used in the initializer as well as the secondary set, we see that for $\varepsilon=\tilde{\Omega}(n^{-1/2})$ and a constant $c$ independent of $\varepsilon$ and $n$ that the total number of vertices used is at most $n$, which means our construction has the right number of vertices.
 How many bad edges does our construction yield?
-Starting from the first layer in the secondary set, we can turn each successive layer red in sequence so that the last two layers form a $K_{c\cdot\varepsilon n,c\cdot\varepsilon n}$ where the left side is red and the right half is white.
+Starting from the first layer in the secondary set, we can turn each successive layer red in sequence so that the last two layers form a $K_{c\cdot\varepsilon n,c\cdot\varepsilon n}$ where the left side is red and the right half is blue.
 However, this produces only $c^{2}\cdot\varepsilon^{2}n^{2}$ bad edges, which, recalling that we started with $\frac{1}{\varepsilon}$ bad edges, matches the bound $\Omega(\varepsilon^{3}n^{2})$ on the price of uncertainty in Balcan et al.
@@
 It turns out we can produce many more bad edges in our construction by doing the following.
-Note that the last layer in our construction is white, and all layers between the frozen layers defined above and the last layer are now red.
-We now turn each vertex in the layer after the frozen layers white, which is a valid uncertain best response by Proposition 3.
-We can now turn each successive layer white in sequence until the last three layers are white, red, and white respectively.
-Repeatedly switching the colors of the vertices in the layer after the frozen layers and propagating uncertain best responses, we can reach a state in which the boosting operation layers alternate between white and red.
-Every edge between boosting operation layers is now bad, which makes $\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{1-\left(\frac{1}{1+\frac{\varepsilon}{2}}\right)^{2}}\right)=\Theta(\varepsilon n^{2})$ bad edges at the ending state.
-Recalling that we began with $\frac{1}{\varepsilon}$ bad edges, we have established that the price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ in the domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
+Note that the last layer in our construction is blue, and all layers between the frozen layers defined above and the last layer are now red.
+We now turn each vertex in the layer after the frozen layers blue, which is a valid uncertain best response by Proposition 3.
+We can now turn each successive layer blue in sequence until the last three layers are blue, red, and blue, respectively.
+By repeatedly switching the colors of the vertices in the layer after the frozen layers and propagating uncertain best responses, we can reach a state in which the boosting operation layers alternate between blue and red.
+Every edge between boosting operation layers is now bad, which makes
+$\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{1-\left(\frac{1}{1+\frac{\varepsilon}{2}}\right)^{2}}\right)=\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{\frac{\varepsilon(4+\varepsilon)}{(2+\varepsilon)^{2}}}\right)=\Theta(\varepsilon n^{2})$
+bad edges at the ending state.
+Recalling that we began with $\frac{1}{\varepsilon}$ bad edges, we have established that the price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ in the domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
 4 Upper Bound
@@
 The PoU for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$.
-In addition to focusing only on the steps at which each vertex plays its first uncertain best response, we will also carefully construct and analyze a monovariant that restricts the degrees of new vertices and allows us to bound the number of edges in the graph.
+In addition to focusing only on the steps at which each vertex plays its first uncertain best response, we will also carefully construct and analyze a construction that restricts the degrees of new vertices and allows us to bound the number of edges in the graph.
 4.1 Proof of Improved Upper Bound
 Let $G$ be a graph for a consensus game, and let $V$ be the set of vertices of $G$. It will be useful to first establish a bound upon the starting number of bad edges.
-Proposition 5.
+Claim 5.
 Suppose that it is possible to increase the number of bad edges in $G$ by a factor of $\Omega\left(\varepsilon^{2}n^{2}\right)$.
@@
 Then, for all integers $1\leq k\leq m$, define $S_{k}=S_{0}\cup\{p_{1},p_{2},\dots,p_{k}\}$.
+Notice that the first vertex $p_{1}$ can only change its color after several vertices in $S_{0}$ change colors, because $p_{1}$ must begin with only good edges.
 Now that we have set up our notation, we will prove some preliminary statements regarding the behavior of the consensus game.
@@
 Proof.
-Clearly, $m=|P|<|V|=n$, so $m=O(n)$ and the upper bound holds. We will now attack the lower bound.
+Clearly, $m=|P|<|V|=n$, so $m=O(n)$ and the upper bound holds. We will now prove the lower bound.
 By construction, $V$ consists of the two disjoint subsets $S_{m}$ and $Q$.
@@
 This means that we can delete these vertices along with all of their incident edges, which will never decrease the factor by which the number of bad edges in $G$ can be increased.
-By construction, prior to when $p_{1}$ plays its first best response, any edges between 2 vertices in $V\setminus S_{0}$ must be good.
-Since $Q=V\setminus S_{m}\subset V\setminus S_{0}$, and since vertices in $Q$ never change their colors, any edges between two vertices in $Q$ must always be good.
-By removing all such edges, we will never decrease the factor by which the number of bad edges in $G$ can be increased.
-We are now left with a new set of vertices $Q^{\prime}\subset Q$ where every edge incident to a vertex in $Q^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=P$.
-For integers $1\leq k\leq m$, $p_{k}$ can only play its first best response if the number of its bad neighbors is at least the number of its good neighbors multiplied by $(1+\varepsilon)^{-1}$.
-By Proposition 6, the number of bad edges incident to $p_{k}$ is at most $|S_{k-1}|<|S_{m}|$.
-Right before $p_{k}$ plays its first best response, all edges between $p_{k}$ and a vertex in $Q^{\prime}$ must be good, so
-$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|Q^{\prime}|$
-$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|Q^{\prime}|$
-$\Theta\left(|S_{m}|\right)$ $=\Theta\left(Q^{\prime}\right).$
-Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$.
-Recall that $|S_{m}|=|S_{0}|+m$.
-By Proposition 5, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin with a bad neighbor is at most twice this quantity.
-So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.
-$\Theta(n)$ $=|S_{m}|$
-$\Theta(n)$ $=|S_{0}|+m$
-$\Theta(n)$ $=O\left(\varepsilon^{-2}\right)+m$
+We will take this a step further by deleting all vertices in $Q$ that do not share an edge with a vertex in $P$.
+By similar logic, this cannot decrease the PoU because all vertices in $V\setminus Q$ can still have at most as many good edges as before.
+Thus, the same sequence of uncertain best responses can be played to produce the same number of final bad edges even though we now have fewer vertices in $V$.
+After all of our deletions, we are now left with a new set of vertices $Q^{\prime}\subset Q$ where every edge incident to a vertex in $Q^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=P$.
+Then, there exists a vertex $v\in P$ with the most edges connecting it to a vertex in $Q^{\prime}$.
+Let the number of such edges be $w$.
+We then construct a set $R$ of vertices $\{r_{1},r_{2},\dots,r_{w}\}$, and we can add new edges such that each vertex in $P$ connected by $j$ edges to vertices in $Q^{\prime}$ is connected to $r_{1},r_{2},\dots,r_{j}$.
+Finally, we delete all vertices in $Q^{\prime}$ and all of their incident edges.
+The factor by which the number bad edges in $G$ can be increased doesn’t change because all vertices that change colors still have the same number of good edges.
+Furthermore, $|R|=w\leq|Q^{\prime}|$, so the PoU cannot decrease.
+Finally, since $v$ always has at least $w=|R|$ good edges, it must have at least $|R|(1+\varepsilon)^{-1}=\Theta(|R|)$ bad edges when it plays its first uncertain best response.
+However, by construction, $v$ cannot have more than $|S_{m}|$ bad edges, so $|R|=O(|S_{m}|)$.
+Since the entire graph now only consists of vertices in $R$ and $S_{m}$, this shows that $n=\Theta(|S_{m}|)$.
+Next, recall that $|S_{m}|=|S_{0}|+m$.
+By 5, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin with a bad neighbor is at most twice this quantity.
+So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well, and we thus have
+$\Theta(n)$ $=|S_{m}|=|S_{0}|+m=O\left(\varepsilon^{-2}\right)+m.$
 Since $\varepsilon=\omega\left(n^{-1/2}\right)$, this becomes
-$\Theta(n)$ $=o\left(n\right)+m$
-$\Theta(n)$ $=m.$
+$\Theta(n)$ $=o\left(n\right)+m\implies\Theta(n)=m.$
 With our construction, Proposition 6 allows us to contain the bad edges in the graph at any step inside one of $S_{0},S_{1},\dots,S_{m}$.
@@
 Example 9.
-Suppose that $A_{0}=\{3,2,2\}$, $\varepsilon=1/4$, and we wish to play a move on $z=4$. Then, $\lfloor(1+\varepsilon)z\rfloor=5$, and the possible multisets for $A_{1}$ are:
-$\{3,3,1,1,1\},\quad\{3,2,1,1,1,1\},\;\;\text{and}\;\;\{2,2,1,1,1,1,1\}.$
+Suppose that $A_{0}=\{3,2,2\}$, $\varepsilon=1/3$, and we wish to play a move on $z=3$. Then, $\lfloor(1+\varepsilon)z\rfloor=4$, and the possible multisets for $A_{1}$ are:
+$\{3,3,1,1\},\quad\{3,2,1,1,1\},\;\;\text{and}\;\;\{2,2,1,1,1,1\}.$
 The new multiset $A_{k}$ represents the edges between $p_{k}$ and a vertex in $V\setminus S_{k}$.
@@
   • Delete one instance of $z$ from $A_{k}$.
-  • Pad $A_{k}$ with an sufficiently large (at least $z(1+\varepsilon)$) number of zeroes.
+  • Pad $A_{k}$ with an sufficiently large (at least $\lfloor z(1+\varepsilon)\rfloor$) number of zeroes.
   • Create $A_{k+1}$ from $A_{k}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $A_{k}$ and adding 1 to each of them.
-  • Remove all zeroes from $A_{k+1}$.
+  • Remove all remaining zeroes from $A_{k+1}$.
 As a shorthand notation, let $\sum A_{k}$ denote the sum of the elements (with their multiplicities) of $A_{k}$, and let $\sum A_{k}^{2}$ denote the sum of squares of the elements (with their multiplicities) of $A_{k}$.
@@
 Let $T_{k}$ denote the number of edges between two vertices in $S_{k}$.
-Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum A_{k}$ denote $(\sum A_{k})-(\sum A_{k-1})$.
 For every regular move that we play on $A_{k-1}$ with an integer $z\in A_{k-1}$, the sum $\sum A_{k}$ grows from $\sum A_{k-1}$ by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges between two vertices in $S_{k}$ grows from $S_{k-1}$ by $z$.
 This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$.
-We have
-$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum A_{k}$
-$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum A_{k}$
-$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum A_{k}$
+Thus, we have
+$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\sum A_{k}-\sum A_{k-1},$
+as well as
+$\varepsilon\left(T_{k}-T_{k-1}\right)-1=\varepsilon z-1<\lfloor\varepsilon z\rfloor,$
+which by substitution implies the inequality
+$\varepsilon\left(T_{k}-T_{k-1}\right)-1<\sum A_{k}-\sum A_{k-1}.$
 By summing both sides from $k=1$ to $m$, we get
-$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum A_{m}-\sum A_{0}.$
+$\varepsilon\left(T_{m}-T_{0}\right)-m$ $\leq\sum A_{m}-\sum A_{0}.$
 Since $\sum A_{m}=O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$,
-$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}.$
+$\varepsilon\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}.$
 By Lemma 7, this becomes
-$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}$
+$\varepsilon\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}$
 $\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}+\Theta(n)+\varepsilon T_{0}$
@@
 By Proposition 11, $\varepsilon^{-1}\sum A_{0}$ must be $O(\varepsilon^{-1}n|S_{0}|)$.
-Since $\varepsilon=\omega(n^{-1/3})$, this is dominated by the $O(\varepsilon^{2}n^{2}|S_{0}|)$ term.
+Since we assumed that $\varepsilon=\omega(n^{-1/3})$, this is dominated by the $O(\varepsilon^{2}n^{2}|S_{0}|)$ term.
 Also, since we must have $\Omega(\varepsilon^{-1})$ starting bad edges, $S_{0}$ is $\Omega(\varepsilon^{-1})$ as well.
-This means that $O(\varepsilon n^{2})$ must also be $O(\varepsilon^{2}n^{2}|S_{0}|)$.
-Finally, $\varepsilon=\omega(n^{-1/3})$, so $O(\varepsilon n^{2})$ will dominate $\Theta(\varepsilon^{-1}n)$.
+This means that $O(\varepsilon n^{2})$ must also be $O(\varepsilon^{2}n^{2}|S_{0}|)$, and then $\varepsilon=\omega(n^{-1/3})$ implies that $O(\varepsilon n^{2})$ will dominate $\Theta(\varepsilon^{-1}n)$.
 We are left with the following terms on the RHS:
@@
 By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$.
-By Proposition 5, we start with $O(\varepsilon^{-2})$ bad edges, so $T_{0}=O(\varepsilon^{-2})$.
+By 5, we start with $O(\varepsilon^{-2})$ bad edges, so $T_{0}=O(\varepsilon^{-2})$.
 This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega(n^{-1/3})$.
@@
 Having the sum $\sum A_{k}$ spread out across many small elements will make increasing $\sum A_{k}$ in future moves more difficult.
-We consider $\sum A_{k}^{2}$ as a way to measure how ”spread out” the multiset $A_{k}$ is, and this particular quantity is preferable over other measures because the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ can be related to $\sum A_{k-1}$.
+Recall that $\sum A_{k}^{2}$ denotes the sum of squares of the not necessarily distinct elements of $A_{k}$.
+We consider $\sum A_{k}^{2}$ as a way to measure how ”spread out” the multiset $A_{k}$ is, and this particular quantity is preferred over other measures because the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ can be related to $\sum A_{k-1}$.
 Our plan is to bound the behavior of $\sum A_{k}$ and $\sum A_{k}^{2}$ as $k$ increases. The multisets $A_{0}$ and $A_{m}$ will be subject to the following inequalities:
@@
 Eq.
-1 is true because the elements of $E_{m}$ must be nonnegative integers, and Eq.
-2 follows from the Cauchy-Schwarz inequality.
+1 is true because the elements of $A_{m}$ must be nonnegative integers, and Eq.
+2 follows from expanding both sides.
 We will only consider these inequalities (and not those with $m$ and 0 swapped) because we only need to prevent too many regular moves from being played with large $z$, which will naturally cause $\sum A_{k}^{2}$ to become small as $k$ increases.
@@
 Then, pick some element $z\in A_{k-1}$, and define $\zeta=\lfloor z(1+\varepsilon)\rfloor$.
-We can play a regular move on $A_{k-1}$ with $z$, which will involve incrementing $\zeta$ the elements $a_{1},a_{2},\dots,a_{\zeta}\in A_{k-1}$ to $a_{1}+1,a_{2}+1,\dots,a_{\zeta}+1\in A_{k}$.
+We can play a regular move on $A_{k-1}$ with $z$, which will involve incrementing the $\zeta$ elements $a_{1},a_{2},\dots,a_{\zeta}\in A_{k-1}$ to $a_{1}+1,a_{2}+1,\dots,a_{\zeta}+1\in A_{k}$.
 It is easy to see that the effect of such a move on the increase from $\sum A_{k-1}$ to $\sum A_{k}$ is
@@
 Next, we must have
-$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+\sum_{i=1}^{\zeta}\left((a_{i}+1)^{2}+a_{i}^{2}\right)$
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+\sum_{i=1}^{\zeta}\left((a_{i}+1)^{2}-a_{i}^{2}\right)$
 $\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta.$ (4)
@@
 However, we can establish bounds upon these equations that are easier to analyze.
-Since $a_{1},a_{2},\dots,a_{\alpha}$ are all elements of $A_{k-1}$, their sum must be bounded by $\sum A_{k-1}$.
+Since $\varepsilon<1$, we have $\zeta=\lfloor z(1+\varepsilon)\rfloor\leq 2z$, so we can rewrite Eq.
+4 as
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum_{i=1}^{\zeta}a_{i}-z^{2}+2z.$
+Next, since $a_{1},a_{2},\dots,a_{\alpha}$ are all elements of $A_{k-1}$, their sum must be bounded by $\sum A_{k-1}$.
 By plugging this into our expression for the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$, we get
-$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta$
-$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum_{i=1}^{\zeta}a_{j}-z^{2}$
-$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum A_{k-1}-z^{2}.$ (5)
+$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum A_{k-1}-z^{2}+2z.$ (5)
 Proposition 13.
@@
 1.
-This means that always taking the upper bound for $\sum A_{k}$ from Eq.
+This means that always taking our upper bound on $\sum A_{k}$ from Eq.
 5 will never decrease $\sum A_{m}$.
@@
 For the rest of the proof, we will simply assume that $z$ is a positive real number.
-The next step is to find an upper bound for the change from $\sum A_{k-1}$ to $\sum A_{k}$.
+The next step is to find an upper bound on the change from $\sum A_{k-1}$ to $\sum A_{k}$, which is
 $\sum A_{k}-\sum A_{k-1}$ $=\zeta-z$
@@
 The RHS of this equation is clearly positive, so $z^{\prime}$ must be nonpositive as well. We can also bound the size of $z^{\prime}$ with
-$\varepsilon z^{\prime}$ $=\lfloor z(1+\varepsilon)\rfloor-z$
-$\varepsilon z^{\prime}$ $\leq z(1+\varepsilon)-z$
-$z^{\prime}$ $\leq z.$
-By (5), smaller values of $z$ yield larger values of $\sum A_{k}^{2}$. This is exactly what we wanted. ∎
+$\varepsilon z^{\prime}$ $=\lfloor z(1+\varepsilon)\rfloor-z\leq z(1+\varepsilon)-z=\varepsilon z.$
+Thus, $z^{\prime}\leq z$. By Eq. 5, smaller values of $z$ yield larger values of $\sum A_{k}^{2}$. This is exactly what we wanted. ∎
 With Proposition 13 and Proposition 14, we now have a strictly stronger set of actions on $\sum A_{k-1}$ and $\sum A_{k-1}^{2}$ that we can instead use with each of our moves.
@@
 $\sum A_{k}$ $=\sum A_{k-1}+\varepsilon z+1$
-$\sum A_{k}^{2}$ $=\sum A_{k-1}^{2}+2\sum A_{k-1}-z^{2}.$
+$\sum A_{k}^{2}$ $=\sum A_{k-1}^{2}+2\sum A_{k-1}-z^{2}+2z.$
 For all integers $1\leq k\leq m$, let $z_{k}$ denote the value of $z$ used in the strong reversed move from $(\sum A_{k-1},\sum A_{k-1}^{2})$ to $(\sum A_{k},\sum A_{k}^{2})$. Then,
@@
 The same process of expanding is more complicated for $\sum A_{m}^{2}$:
-$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{i}\right)-\sum_{i=1}^{m}z_{i}^{2}.$
-We can rewrite the $\sum A_{j}$ by plugging in Eq. 1 and Eq. 7.
-$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{0}+i+\varepsilon\sum_{j=1}^{i}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
-$0$ $\leq\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\varepsilon\sum_{j=1}^{m}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
-$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum_{i=1}^{m}z_{i}^{2}$ (8)
+$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{i}\right)-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}.$
+We can then rewrite the $\sum A_{j}$ by plugging in Eq. 1 and Eq. 7, and we thus have
+$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{0}+i+\varepsilon\sum_{j=1}^{i}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}$
+$0$ $\leq\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\varepsilon\sum_{j=1}^{m}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}$
+$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}.$ (8)
 By the Cauchy-Schwarz Inequality,
@@
 $\sum_{j=1}^{m}z_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}.$
-Plugging this result into Eq. 8 yields a quadratic polynomial in terms of $\sum_{i=1}^{m}z_{i}$ in the RHS.
-$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
-$0$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}-2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum A_{0}^{2}$
-Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}z_{j}$ must be less than the larger root of this quadratic.
-This larger root is obtained through a tedious application of the quadratic formula:
-$\sum_{j=1}^{m}z_{j}\leq\frac{2\varepsilon m+\sqrt{4\varepsilon^{2}m^{2}+4m^{-1}\sum A_{0}^{2}}}{2m^{-1}}$ (9)
+Plugging this result into Eq. 8 yields the following quadratic polynomial in terms of $\sum_{i=1}^{m}z_{i}$ in the RHS:
+$0$ $\leq\sum A_{0}^{2}+(2\varepsilon m+2)\sum_{i=1}^{m}z_{i}-m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
+$0$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}-(2\varepsilon m+2)\sum_{i=1}^{m}z_{i}-\sum A_{0}^{2}.$
+Since the leading coefficient is positive and the quadratic polynomial is on the lesser side of the inequality, $\sum_{j=1}^{m}z_{j}$ must be less than the larger root of this polynomial.
+This larger root is obtained via the quadratic formula:
+$\sum_{j=1}^{m}z_{j}\leq\frac{(2\varepsilon m+2)+\sqrt{(2\varepsilon m+2)^{2}+4m^{-1}\sum A_{0}^{2}}}{2m^{-1}}.$ (9)
 Lemma 16.
-Suppose that $\varepsilon=\Omega(n^{-1/4})$. Then, $\sum_{i=1}^{m}z_{j}$ is $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
+Suppose that $\varepsilon=\Omega(n^{-1/4})$. Then, $\sum_{i=1}^{m}z_{j}$ must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$.
 Proof.
-If we substitute asymptotic behavior into the RHS of (9), we get
+Since $\varepsilon=\Omega(n^{-1/4})=\Omega(m^{-1/4})$ by Lemma 7, the $2\varepsilon m$ in Eq. 9 must be $\omega(1)$. Thus, if we substitute asymptotic behavior into the RHS of Eq. 9, we get
 $\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum A_{0}^{2}\right)}}{\Theta\left(m^{-1}\right)}.$
@@
 In this case, we can rewrite Eq. 10 as follows:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)}}{\Theta\left(n^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)}{\Theta\left(n^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon n^{2}\right).$
+$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\Theta(\varepsilon n)}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon n^{2}\right).$
 Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$ because we must begin with that many bad edges.
-This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
+This means that $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ must include $O\left(\varepsilon n^{2}\right)$, and we are done with this case.
 Case 16.2.
-$\Theta\left(m|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}m^{2}\right)$.
+$\Theta\left(n|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}n^{2}\right)$.
 In this case, we can rewrite Eq. 10 as follows:
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(n^{3/2}|S_{0}|\right).$
+$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
+$\sum_{j=1}^{m}z_{j}$ $=O\left(n^{3/2}|S_{0}|\right).$
 Since $\varepsilon=\Omega(n^{-1/4})$,
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}n^{3/2}|S_{0}|\right)$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(n^{-1/4}\right)^{-2}n^{3/2}|S_{0}|\right)$
-$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$
-The result of Lemma 16 can be plugged into Eq. 7 with Proposition 11 to obtain a bound on $\sum E_{m}$. Since $\varepsilon=\Omega(n^{-1/4})$, we have
-$\sum A_{m}$ $=\sum A_{0}+\varepsilon\sum_{i=1}^{m}z_{i}$
-$\sum A_{m}$ $=O\left(n|S_{0}|\right)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$
+$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}n^{3/2}|S_{0}|\right)$
+$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon^{2}\left(n^{-1/4}\right)^{-2}n^{3/2}|S_{0}|\right)$
+$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right).$
+The result of Lemma 16 can be plugged into Eq. 7 with Lemma 7 and Proposition 11 to obtain a bound on $\sum A_{m}$. Since $\varepsilon=\Omega(n^{-1/4})$, we have
+$\sum A_{m}$ $=\sum A_{0}+m+\varepsilon\sum_{i=1}^{m}z_{i}$
+$\sum A_{m}$ $=O\left(n|S_{0}|\right)+\Theta(n)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$
 $\sum A_{m}$ $=O\left(\varepsilon^{3}n^{2}|S_{0}|\right).$ (11)